### v2

# It’s High Time : A Survey of Temporal Information Retrieval and Question Answering

Abstract
Time plays a critical role in how information is generated, retrieved, and interpreted. In this survey, we provide a comprehensive overview of Temporal Information Retrieval and Temporal Question Answering, two research areas aimed at handling and understanding time-sensitive information. As the amount of time-stamped content from sources like news articles, web archives, and knowledge bases increases, systems must address challenges such as detecting temporal intent, normalizing time expressions, ordering events, and reasoning over evolving or ambiguous facts. These challenges are critical across many dynamic and time-sensitive domains, from news and encyclopedias to science, history, and social media. We review both traditional approaches and modern neural methods, including those that use transformer models and Large Language Models (LLMs). We also review recent advances in temporal language modeling, multi-hop reasoning, and retrieval-augmented generation (RAG), alongside benchmark datasets and evaluation strategies that test temporal robustness, recency awareness, and generalization.
It’s High Time
: A Survey of Temporal Information Retrieval and Question Answering
1 Introduction
From analyzing centuries-old texts, understanding historical events, to answering questions about emerging developments, time shapes how we seek and interpret information. As digital content continues to grow exponentially across time-stamped sources like news archives, social media, and knowledge bases, the ability to process and reason over temporal information has become essential Alonso et al. (2007). Temporal IR, which searches time-stamped documents, and Temporal QA, which answers time-sensitive queries, together address these needs. Both, collectively referred as Temporal IR/QA, aim to incorporate time-awareness to adapt results to specific periods and resolve time-sensitive queries Campos et al. (2014).
Temporal IR/QA faces distinct challenges that set it apart from standard IR/QA settings. These include identifying temporal intent in queries, interpreting expressions such as "post-World War II" or "in 1998," and modeling relationships between events and their timelines Berberich et al. (2010). Queries may target past, present, or future events, and require systems to identify relevant time frames, order events, and resolve implicit temporal cues. Overcoming these obstacles demands methods that extend beyond traditional keyword-based search and basic retrieval techniques.
For Example, in Figure 1, Q1: "At what age did Obama win the Nobel Peace Prize?" requires constructing a chronology of events by identifying and grounding two temporal anchors, Obama’s birth year (1961) and the year he received the Nobel Peace Prize (2009). The model must then establish a temporal relationship and apply reasoning to compute the answer: 48 years old. Q2: “What does President Obama’s climate policy tell us about how the U.S. viewed climate change during his late years of service?” demands understanding of the query’s intended time and contextual temporal grounding. The model must recognize relative temporal expressions such as “today, next week” and associate them with a reference time such as the document’s publication date. It also needs to retrieve or reason over documents written during the relevant policy timeframe, reconstructing the contemporaneous narrative.
Research in Temporal IR/QA has evolved significantly, building on early foundations to address increasingly complex temporal challenges. Initial efforts relied on rule-based systems Harabagiu and Bejan (2005) and statistical models Berberich et al. (2010) that used document timestamps and hand-crafted rules to interpret time-related information Li and Croft (2003). While these methods established key principles, they struggled to scale or handle diverse temporal contexts. The rise of pre-trained language models has transformed the field by enabling robust temporal reasoning Jain et al. (2023), event sequencing Lin et al. (2021), and adaptation to evolving knowledge Han et al. (2021). These advancements paved the way for more dynamic and scalable temporal systems.
While prior surveys111For a discussion of previous related surveys, we refer the reader to Section A.1. have explored general IR/QA methods Robertson et al. (2009); Xiong et al. (2020); Formal et al. (2021), or focused narrowly on specific aspects of temporal processing within one of these fields Kobayashi and Takeda (2000); Kanhabua et al. (2015); Campos et al. (2014), a comprehensive and unified overview of Temporal IR/QA is long overdue. The most recent dedicated survey in this area was published nearly a decade ago Campos et al. (2014). Since then, the field has grown substantially, driven by advances in Language Models, new datasets, and complex temporal tasks. A survey that captures recent advancements and outlines future research directions is then essential to foster progress and guide the community in developing time-aware systems. Our paper addresses this gap. We trace the evolution from traditional to neural approaches, highlight advances in tasks such as event dating, temporal modeling, and knowledge updating, and outline emerging challenges. In Figure 2, we portray a taxonomy overviewing temporal tasks, datasets, and approaches we will discuss in our review.
2 Key Concepts
We first introduce the core concepts related to Temporal IR/QA.
Temporal Information Retrieval (TIR) aims to retrieve documents that are not only topically relevant but also aligned with the query’s temporal intent. Temporal intent may be explicit such as “Olympics 2024”, or implicit ones, such as “latest Apple earnings”. TIR relies on different temporal signals such as document timestamps (publication dates), temporal expressions (“March 2023”), and event mentions (“2024 Olympics”) to assess a document’s temporal relevance indicating how well its temporal scope matches the query Kanhabua and Nørvåg (2008); Singh et al. (2016).
Temporal Question Answering (TQA) focuses on answering questions with temporal constraints, either explicitly stated, such as "Who won the Nobel Prize in Physics in 2020?" or implied, for instance, "What are the latest US climate policies?". Success in TQA requires understanding the question’s temporal intent and retrieving documents relevant to the corresponding time frame or ones published around that time.
Temporal IR/QA rely on diverse temporal elements. Temporal signals are, in general, defined as features that convey time-related information in text. These include explicit temporal expressions like "March 2023" (used for indexing and filtering), implicit cues such as "recently" (requiring contextual interpretation), relative expressions like "last week" (necessitating to be anchored to a reference point), event-based references such as "2024 Olympics" (linking to known event timelines), and temporal metadata such as document timestamps which indicate publication time and often serve as proxies for judging freshness of content.
The concept of document focus time Jatowt et al. (2013) is crucial here. It denotes the specific time point or interval a document relates to. For example, a 2013’s publication discussing the 2010 Academy Awards has a focus time of 2010 while having a timestamp of 2013. Accurate focus time estimation of documents, using techniques like burst detection, temporal expression analysis, or timestamping named entities, enhances answer precision, especially in news or historical corpora Wang et al. (2020).
forked edges,
for tree=
grow=east,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=left,
font=,
rectangle,
rounded corners,
align=left,
minimum width=0.1em,
edge+=darkgray, line width=0.8pt,
s sep=2pt,
inner xsep=2pt,
inner ysep=2pt,
draw=hidden-draw,
ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center,
where level=1text width=4.7em,font=,,
where level=2text width=5em,font=,,
where level=3text width=6.8em,font=,,
where level=4text width=10em,font=,,
where level=5text width=6.8em,font=,,
Temporal IR/QA, draw=black, fill=gray!10, thick, text=black, ver
Dataset &
Benchmarks ( §3),
color=orange!90, fill=orange!20, thick, text=black
Temporal Document
Collections ( §3.1),
color=orange!90, fill=orange!20, thick, text=black
TQA Datasets ( §3.2),
color=orange!90, fill=orange!20, thick, text=black
TIR Dataset ( §3.3),
color=orange!90, fill=orange!20, thick, text=black
Temporal
Tasks ( §4),
color=cyan!100, fill=cyan!15, thick, text=black
Prediction,
color=cyan!100, fill=cyan!15, thick, text=black
Approaches ( §5),
color=red!80, fill=red!15, thick, text=black
Rule-Based &
Statistical Models ( §5.1),
color=red!80, fill=red!15, thick, text=black
Temporal Language
Models ( §5.2),
color=red!80, fill=red!15, thick, text=black
Temporal RAG ( §5.2.1),
color=red!80, fill=red!15, thick, text=black
Temporal Reasoning
( §5.2.2),
color=red!80, fill=red!15, thick, text=black
In Appendix B we discuss other related concepts, such as temporal taggers, temponyms, granularity, temporal reasoning, timeline extraction, disambiguation, and robustness.
Dataset #Questions Knowledge Source Creation Method Answer Type Time Frame Temporal Metadata Multi-Hop NewsQA Trischler et al. (2017) 119k News CS Freeform 2007-2015 ✗ ✗ TDDiscourse Naik et al. (2019) 6.1k News CS Extractive Unspecified ✗ ✗ TORQUE Ning et al. (2020) 21k News CS Abstractive - ✗ ✗ ArchivalQA Wang et al. (2022) 532k News AG Extractive 1987-2007 ✓ ✗ TimeQA Chen et al. (2021) 41.2K Wikipedia AG Extractive 1367-2018 ✗ ✗ TiQ Jia et al. (2024) 10K Wikipedia AG Freebase Unspecified ✗ ✗ TempQuestions Jia et al. (2018) 1.2k Freebase AG Extractive Unspecified ✗ ✓ TemporalQuestions Wang et al. (2021a) 1K News CS Extrcative 1987-2007 ✓ ✗ TempLAMA Dhingra et al. (2022) 50k News CS Extractive 2010-2020 ✓ ✗ ComplexTempQA Gruber et al. (2024) 100,228k Wikipedia AG Extractive 1987-2023 ✓ ✓ MenatQA Wei et al. (2023) 2.8k Wikipedia AG Extractive 1367-2018 ✗ ✗ PAT-Question Meem et al. (2024) 6,1k Wikipedia CS Extractive - ✗ ✓ TempTabQA Gupta et al. (2023) 11.4k Wikipedia Info box CS Abstractive - ✗ ✗ SituatedQA Zhang and Choi (2021) 12.2k Wikipedia CS – ✗ ✗ UnSeenTimeQA Uddin et al. (2024) 3.6k Synthetic AG Abstractive - ✗ ✓ ChroniclingAmericaQA Piryani et al. (2024b) 485k News AG Extractive 1800-1920 ✓ ✗ FRESHQA Vu et al. (2024) 600 Google Search CS - - ✗ ✓ COTEMPQA Su et al. (2024) 4.7k Wikidata CS Abstractive ✗ ✓ Test of Time (ToT) Fatemi et al. (2024) 1.8k Synthetic AG Abstractive - ✗ ✓ TIMEDAIL Qin et al. (2021) 1.1k DailyDialog CS Multiple-choice - ✗ ✗ Complex-TR Tan et al. (2024) 10.8 Wikipedia+Google Search AG Multi-answer ✗ ✓ StreamingQA Liska et al. (2022) 147k News CS Extractive 2007-2020 ✓ ✓ TRACIE Zhou et al. (2021) 5.4k Wikipedia CS abstractive ✗ ✗ ForecastQA Jin et al. (2021) 10.3k News CS Multiple-Choice 2015-2019 ✓ ✓ TEMPREASON Tan et al. (2023) 52.8k Wikipedia/Wikidata SC Abstractive 634-2023 ✗ ✗ TemporalAlignmentQA Zhao et al. (2024) 20k Wikipedia AG Abstractive 2000-2023 ✗ ✗ ReaLTimeQA Kasai et al. (2023) 5.1k Search CS Multiple-choice 2020-2024 ✗ ✗
3 Datasets and Evaluation Benchmarks
The development of Temporal IR/QA systems fundamentally depends on the availability of temporally grounded datasets and robust evaluation methodologies used for training, testing, and benchmarking time-aware models. We provide an overview in this section of temporal datasets, organized into three categories: temporal document collections, TQA datasets, and TIR datasets.
3.1 Temporal Document Collection
Prior work has utilized diachronic and synchronic document collections as well as annotated temporal corpora.
Diachronic corpora consist of time-stamped documents spanning extensive time periods. They support retrospective retrieval, diachronic analysis, and event-based reasoning. Prominent examples include the New York Times Annotated Corpus (1987–2007; 1.8m articles) Sandhaus (2008), which, for example, serves as the basis for ArchivalQA Wang et al. (2022) dataset, and the CNN/Daily Mail corpus (2007–2015; 313k articles) Hermann et al. (2015) used, e.g., in NewsQA Trischler et al. (2017). The Chronicling America collection (1800–1920) offers digitized historical newspaper articles and supports long-range historical QA via ChroniclingAmericaQA Piryani et al. (2024b). More recently, the Newswire corpus Silcock et al. (2024) has expanded the length of time frames, providing 2.7 million newswire articles published between 1878 and 1977. It is enriched with metadata including geo-referenced datelines, Wikipedia/Wikidata entity links, and topical annotations, enabling fine-grained historical and spatio-temporal modeling. Another widely used corpus is CUSTOMNEWS Lazaridou et al. (2021) (1969–2019), which consists of crawled English news sources and spans diverse domains including politics, finance, and sports.
Diachronic corpora are also used in a range of related temporal tasks, including semantic drift detection Hamilton et al. (2016), event burst modeling Radinsky and Horvitz (2013), and timeline construction Gutehrlé et al. (2022).
Synchronous corpora represent a coherent snapshot of the world at a specific point in time. Unlike diachronic corpora, which typically span decades or years, synchronous collections capture a temporally aligned view, sometimes in conjunction with structured knowledge bases. Wikipedia articles Vrandečić and Krötzsch (2014), for example, reflect a particular version of world knowledge at a certain time (when the dump was made) and can be linked to Wikidata timestamps. Datasets like TimeQA Chen et al. (2021), TEMPREASON Tan et al. (2023), and ComplexTempQA Gruber et al. (2024) build on Wikipedia snapshots to support temporally-scoped QA grounded in a time-specific context.
Finally, Annotated temporal corpora with explicit temporal annotations facilitate more structured forms of temporal reasoning. TimeBank Pustejovsky et al. (2003) introduced TimeML to annotate temporal expressions, events, and their temporal relations. Follow-up datasets like WikiWars Mazur and Dale (2010) and RED O’Gorman et al. (2016) extended it to historical narratives and causal relations, respectively. Such corpora constitute gold-standard resources for temporal tagging and relation extraction.
3.2 TQA Datasets
TQA datasets allow evaluating how well systems can answer questions that require temporal reasoning. They vary along multiple dimensions, including Knowledge Source, Temporal Orientation, Temporal Explicitness, and Reasoning Complexity.
Knowledge Source
TQA datasets are commonly derived from diachronic or synchronic corpora. Diachronic Corpora (also known as Primary Sources) tend to provide contemporaneous accounts written around the time when events occurred in the past. Datasets such as NewsQA Trischler et al. (2017), TDDiscourse Naik et al. (2019), TORQUE Ning et al. (2020), ArchivalQA Wang et al. (2022), TKGQA Ong et al. (2023), ChroniclingAmericaQA Piryani et al. (2024b), are curated from old news sources and can be used to evaluate models’ abilities to retrieve and reason over temporally anchored document collections. Table 1 lists all the datasets and the types of knowledge sources used to generate their questions.
In contrast, Synchronic Corpora like Wikipedia basically constitute Secondary Sources since they provide retrospective view of the past. They have been used to build datasets like TimeQA Chen et al. (2021), TEMPREASON Tan et al. (2023), TiQ Jia et al. (2024), and ComplexTempQA Gruber et al. (2024), which support fine-grained reasoning across temporally scoped, consistent knowledge bases.
Recent advancements have also seen the emergence of purely synthetic datasets designed to specifically test models on controlled and complex temporal reasoning scenarios. For example, UnSeenTimeQA Uddin et al. (2024) introduces a novel, data contamination-free benchmark that evaluates temporal reasoning independently from any pre-training knowledge.
Temporal Orientation
While most datasets focus on past events, future-oriented QA datasets remain relatively rare. Still, they are increasingly important for evaluating models’ ability to perform predictive and hypothetical reasoning. ForecastQA Jin et al. (2021) and TimeBench Chu et al. (2024) are among the few benchmarks that include questions about future events, testing models’ ability to perform timeline projections and forecast-based inference.
Question Type
Temporal questions can be broadly classified by their explicitness in referencing time. Datasets like TimeQA Chen et al. (2021), SituatedQA Zhang and Choi (2021) and TempQuestions Jia et al. (2018) contain Explicit Temporal Questions with clear temporal markers, such as "What happened in 1947?", signaling temporal intent directly.
In contrast, Implicit Temporal Questions omit direct time references but still require temporal inference. For instance, "Who was Prime Minister of the UK when the Berlin Wall fell?" requires inferring the date of the event and then linking it to a temporally relevant fact. Datasets such as TiQ Jia et al. (2024) and TORQUE Ning et al. (2020) focus on implicit reasoning, testing event-event and event-time relationships. Others like ArchivalQA Wang et al. (2022), TemporalQuestions Wang et al. (2021a), and ComplexTempQA Gruber et al. (2024) combine both question types, offering a spectrum of temporal reasoning demands from explicit, time-anchored queries to implicit, event-based inference.
Temporal Reasoning Complexity
TQA tasks also vary in the depth of reasoning they require. Simple Temporal Questions typically involve direct lookups, such as identifying the date of a specific event or the state of the world at a given time. Early datasets like NewsQA Trischler et al. (2017) and TempLAMA Dhingra et al. (2022) largely belong to this category. In contrast, Complex Temporal Questions demand more intricate processing such as multi-hop reasoning, temporal filtering, or synthesizing information across events. For example, the question “What major international agreements were signed after World War I but before World War II?” necessitates multi-hop temporal reasoning and contextual comparison. Datasets like MenatQA Wei et al. (2023), TempReason Tan et al. (2023), Complex-TR Tan et al. (2024), and ComplexTempQA Gruber et al. (2024) are explicitly designed to evaluate these advanced reasoning capabilities. Others like TimeBench Chu et al. (2024) span both simple and complex reasoning levels, including tasks such as timeline construction or event duration inference. Table 1 compares various datasets for Temporal QA/IR.
3.3 TIR Datasets
While TQA datasets focus on answering time-sensitive questions, TIR datasets support tasks such as identifying time-sensitive documents, modeling temporal query intent, and ranking documents by temporal relevance or diversity. They typically pair queries with timestamped corpora and are designed to assess retrieval systems’ performance across temporal dimensions.
The Temporalia series at NTCIR-11 and NTCIR-12 Joho et al. (2014, 2016) established foundational benchmarks for TIR through two tasks: Temporal Query Intent Classification (TQIC), which categorizes queries by temporal orientation (e.g., past, recency, future, atemporal), and TIR, which ranks documents based on their temporal relevance or diversity. The tasks use the LivingKnowledge News/Blog Corpus Matthews et al. (2010), containing 3.8 million timestamped documents (2011–-2013) annotated with time expressions and named entities. Apart from Temporalia, TREC Temporal Summarization Track Diaz et al. (2015) offered datasets for a related task of real-time event summarization, testing systems’ ability to rank documents by recency and relevance as well as emphasizing temporal diversity and freshness. In parallel, the TempEval series from the SemEval workshops UzZaman et al. (2013); Verhagen et al. (2010, 2007) provided benchmark datasets for temporal information extraction such as temporal expression, event, and temporal relation, crucial for supporting TIR tasks.
4 Temporal Prediction Tasks
Temporal prediction tasks are essential for developing time-aware IR and QA systems. They focus on inferring implicit or missing temporal information from text, thereby improving the alignment between queries, documents, and events. These tasks are critical when explicit temporal metadata is sparse, noisy, or unavailable, and they support applications such as historical search, timeline construction, and temporally sensitive retrieval.
Key tasks include Event Dating, Document Dating, Focus Time Estimation, Query Time Profiling, and Event Occurrence Prediction. Traditional methods rely on statistical language models and handcrafted rules, while more recent techniques employ transformer-based encoders, temporal embeddings, and graph-based reasoning to improve generalization and robustness Yang et al. (2023); Abdallah et al. (2025); Liu and Quan (2025); Yang et al. (2024). For a detailed review of task definitions, representative techniques, and evaluation strategies, we refer the readers to Appendix C.
5 Approaches in Temporal IR/QA
A wide range of approaches have been developed to address the challenges of Temporal IR/QA, from early rule-based systems and statistical models to neural networks and large language models (LLMs). They differ in how they represent temporal information, reason over temporal relationships, and adapt to changing world knowledge.
5.1 Rule-based & Statistical Methods
Early work in Temporal IR/QA was dominated by rule-based systems and statistical models that laid the groundwork for core temporal tasks such as time expression normalization, event ordering, and temporal ranking. While limited in scalability and adaptability, they introduced many foundational concepts that remain relevant today.
In TIR, rule-based systems focused on extracting and normalizing time expressions to improve retrieval for time-sensitive queries Arikan et al. (2009); Alonso et al. (2007). Models like TCluster Alonso et al. (2009) and time-based language models Li and Croft (2003) used document timestamps and decay functions to model recency, while others like Berberich et al. (2010) combined metadata and vague expressions in probabilistic ranking models. To handle implicit temporal intent, techniques such as median timestamp analysis Kanhabua and Nørvåg (2010) and query log mining Metzler et al. (2009) were introduced.
Other strategies focused on enhancing recency-aware retrieval. Jatowt et al. (2005) proposed re-ranking methods using archived web snapshots to favor fresher content, while Dong et al. (2010) incorporated real-time Twitter signals, and Setty et al. (2017) used news signals into crawling and ranking to support time-sensitive queries. Efficient indexing methods were also developed to support temporal queries over evolving corpora such as Wikipedia and web archives Anand et al. (2011, 2012); Holzmann and Anand (2016). Styskin et al. (2011) introduced a machine learning model to predict recency sensitivity, combining it with greedy diversification to balance freshness and topical relevance.
As TIR matured, researchers began modeling the temporal dynamics of both queries and documents. Kulkarni et al. (2011) analyzed how user intents evolve over time, highlighting the need for adaptive retrieval strategies that can respond to temporal drift in query behavior. Joho et al. (2013) studied the prevalence of different temporal orientations of user queries, and the strategies user apply to find temporally relevant content from the past, future or present. Later systems adapted ranking strategies to temporal query profiles using machine learning Kanhabua et al. (2012) or temporal interval representations Rizzo et al. (2022).
Early QA systems like Harabagiu and Bejan (2005) relied on TimeML and lexical resources like WordNet Miller (1992) for event reasoning. To handle complex temporal questions more effectively, Saquete et al. (2004, 2009) introduced a multi-layered QA architecture that decomposed questions into temporally constrained sub-questions using temporal expression taggers like TERSEO Saquete et al. (2003). These approaches showed improved precision and generalizability across languages.
Despite their simplicity, rule-based and statistical methods introduced key mechanisms of temporal intent modeling, expression normalization, and timeline reasoning that continue to influence more advanced systems.
5.2 Temporal Language Models
The emergence of deep learning has significantly advanced Temporal IR/QA by enabling models to capture temporal dependencies and contextual nuances. Recent research has led to the development of Temporal Language Models (TLMs) that explicitly incorporate temporal signals during pretraining or fine-tuning. Models such as TempoT5 Dhingra et al. (2022), TempoBERT Rosin et al. (2022), and BiTimeBERT Wang et al. (2023) included timestamps and temporal expressions directly into their training inputs or used time-focused pretraining tasks, improving temporal generalization in downstream tasks such as semantic change detection and Temporal QA. Other approaches, like syntax-guided temporal language model (SG-TLM) Su et al. (2023), enhance sensitivity to temporal structure by masking syntactic and semantic spans that carry temporal meaning.
On the other hand, Cao and Wang (2022) explored time-aware generation by introducing temporal prompts, including both natural language timestamp descriptions and continuous vector (linear) representations of timestamps. Beyond input-level integration, time-aware language models like TALM Ren et al. (2023) incorporate time-specific word representations through hierarchical modeling and temporal adaptation, achieving strong results in historical text dating. TCQA Son and Oh (2023) employs synthetic data and a time-context span selection task to train models that align time-aware representations with contextually grounded answers. Further, techniques such as Temporal Span Masking (TSM) Cole et al. (2023) and temporal attention mechanisms Rosin and Radinsky (2022) incorporate explicit temporal annotations into transformer architectures to improve time sensitivity.
5.2.1 Temporal RAG
While TLMs improve temporal understanding through pretraining, they remain limited by the static nature of training data. To address evolving information needs and reduce temporal hallucinations, recent work has turned to Retrieval-Augmented Generation (RAG) that integrates neural retrieval with generation to incorporate up-to-date, time-relevant evidence at inference time.
Recent temporal RAG systems extend this idea by embedding temporal signals directly into retrieval and generation pipelines. TempRALM Gade and Jetcheva (2024) introduces temporal signals into dense retrieval, enhancing recency and factual grounding for time-sensitive queries. TempRetriever Abdallah et al. (2025) and TsContriever Wu et al. (2024) encode temporal relevance directly into dense retrievers, improving alignment between temporal queries and evidence. TimeR4 Qian et al. (2024) proposes a Retrieve-Rewrite-Retrieve-Rerank pipeline that transforms implicit temporal queries into explicit ones, retrieves from time-anchored knowledge sources, and reranks based on temporal constraints. MRAG Siyue et al. (2024) adapts RAG with multi-source and multi-hop temporal retrieval for event-centric QA. To mitigate hallucinations and outdated generations, FRESHPROMPT Vu et al. (2024) integrates real-time signals into the prompting and retrieval process. Together, these models make RAG more responsive to temporal dynamics in IR/QA.
5.2.2 Temporal Reasoning Capabilities
While Temporal Language Models enhance time-aware representations and retrieval, many Temporal IR/QA tasks demand more sophisticated reasoning, such as understanding event sequences, temporal constraints, and durations.
Temporal reasoning capabilities in pre-trained language models (PLMs) have seen notable improvements, with recent efforts focusing on enhancing zero-shot generalization and temporal robustness. Continual temporal adaptation methods, including ECONET Han et al. (2021), enhance temporal relational coherence and consistency across evolving contexts. Structural temporal reasoning models like TIMERS Mathur et al. (2021), and ConTempo Niu et al. (2024) address multi-hop and document-level inference with specialized architectures. Moreover, event duration and ordering prediction have benefited from task-specific temporal pretraining objectives (e.g., E-PRED, R-PRED) Yang et al. (2020) and transfer learning strategies Virgo et al. (2022).
Despite these advancements, modeling temporal relationships in LLMs remains challenging. Recent benchmarks such as TRAM Wang and Zhao (2024) evaluate LLMs on tasks like event ordering, arithmetic, frequency, and duration, revealing that even strong models like GPT-4 fall short of human-level performance. To isolate genuine reasoning abilities from memorization, Test of Time (ToT) Fatemi et al. (2024) introduces synthetic tasks targeting temporal logic and inference. Additionally, TODAY Feng et al. (2023) challenges models with subtle temporal shifts and differential analysis. Methods like Narrative-of-Thought Zhang et al. (2024) guide models to generate structured temporal narratives. Finally, Wallat et al. (2024, 2025) study temporal blind spots of LLMs and their resiliency to changes in time-related elements (e.g., altering a date in a query, or its position) elucidating missing knowledge and showing that current models are still vulnerable to adversarial or other perturbations.
6 Future Directions
Despite two decades of research and significant progress, Temporal IR/QA systems still struggle to adapt to evolving real-world events, shifting user needs, and dynamic data streams. To advance the development of time-aware systems, we propose future directions organized into three core themes: System Design (architectures and real-time capabilities), Knowledge Management (updating and representing time-sensitive knowledge), and Evaluation and Robustness (metrics and generalization). These directions address gaps identified throughout this survey, such as temporal bias, rigidity in models, and the limited scope of existing evaluations.
6.1 System Design
Real-Time Information Integration. Most IR systems depend on periodically updated corpora, leaving them blind to rapidly unfolding events like elections, protests, or trending information. Future work should treat data as a continuous stream, enabling real-time indexing Baeza-Yates and Ribeiro-Neto (2011), burst detection Wang et al. (2021a), and responsive re-ranking Tran et al. (2015), as well as supporting applications like live event tracking or misinformation detection V et al. (2024).
Development of Temporally-Aware LLM Agents.
Current LLM agents prioritize task completion or dialogue but lack structured temporal understanding Wallat et al. (2024). Future systems should include dedicated temporal understanding methods for better understanding temporal references, semantics, and test-time reasoning.
6.2 Knowledge Management
Advanced Temporal Knowledge Editing. Static models struggle to keep up with real-world change. Instead of retraining, future systems could use modular, trackable edit layers for local updates, preserving historical facts.
Integration of Diachronic and Synchronic Knowledge.
Temporal questions often require combining evolving facts (e.g., event timelines) with stable knowledge (e.g., definitions). Future systems should integrate diachronic sources with synchronous sources to provide comprehensive answers. For example, answering "How has the unemployment rate changed since 2008?" requires diachronic trends from datasets like ArchivalQA (Section 3) and synchronous explanations from Wikipedia, addressing the aggregation needs (Section 2).
Multilingual Temporal IR/QA.
Temporal expressions vary across languages and cultures, posing challenges for globalized systems. For instance, date formats differ (e.g., DD/MM/YYYY vs. MM/DD/YYYY), and cultural references (e.g., "post-Meiji era" in Japanese) require context-specific interpretation. Future research should develop cross-lingual temporal taggers, multilingual benchmarks, and culturally adaptive models, building on multilingual taggers like HeidelTime Strötgen and Gertz (2010).
6.3 Evaluation and Robustness
Implicit Temporal Intent Understanding. Many queries imply but do not state a time frame. Future work should improve models’ ability to infer latent temporal scopes using derived labels or event grounding. This addresses the implicit reasoning challenges in datasets like TORQUE Ning et al. (2020) and TiQ Jia et al. (2024).
Robustness to Temporal Drift and Misalignment.
Performance drops when models are applied to data from different time periods, which can reduce accuracy Shin et al. (2025); Zhang and Choi (2023); Luu et al. (2022); Wallat et al. (2025). Future work should enhance model resilience to temporal misalignment, building on the robustness challenges in Test of Time.
7 Conclusion
Temporal IR/QA is critical for retrieving and reasoning over time-sensitive information in dynamic, evolving contexts. In this survey, we have traced the field’s progression from early rule-based systems to TLMs and RAG approaches. We identified core challenges, including temporal tagging, temporal intent detection, event ordering, and robustness to evolving facts and implicit temporal signals.
Our review highlights persistent limitations such as reliance on static knowledge, limited capabilities for future-oriented reasoning, and dataset bias toward past events. We show that temporal complexity, vague expressions, knowledge drift, and real-time demands significantly impact system behavior and evaluation.
Despite notable progress, current systems often struggle with temporal uncertainty, maintaining consistency across time, and adapting to multilingual or culturally diverse temporal expressions. As real-world applications increasingly require temporally adaptive systems, these gaps point to the need for richer evaluation protocols, improved temporal representations, and continual learning strategies. We anticipate future progress toward robust, time-aware IR and QA systems capable of understanding not just what happened, but also when, why, and how information evolves over time.
Limitations
This survey aims to provide a comprehensive overview of Temporal IR/QA. There are a few important limitations to acknowledge.
We made our best efforts to be thorough, but it is possible that some relevant works may have been missed. We conducted an extensive literature review using forward and backward snowballing techniques, with particular attention to papers published in major venues such as ACL, SIGIR, EMNLP, NeurIPS, ECIR, and preprints on arXiv. On the other hand, due to page limitations, we provide only a very brief summary of each method without exhaustive technical details.
Appendix A Appendix
A.1 Related Surveys
Advances in temporal datasets, time-aware models, and reasoning techniques have enabled systems capable of retrieving time-relevant documents, ordering events, and answering temporally constrained questions, benefiting applications such as historical analysis, fact-checking, and intelligent assistants.
While IR and QA have been widely surveyed, most existing reviews focus on general techniques, often neglecting temporal aspects. IR surveys emphasize ranking functions, neural retrieval models, and query understanding Li et al. (2025); Zhu et al. (2023), while QA surveys center on extractive, abstractive, or multi-hop answering over static knowledge sources Zhu et al. (2021); Mavi et al. (2024). They rarely consider temporal intent, dynamic or evolving information, or event sequencing, highlighting a key gap that remains unaddressed.
Several earlier works provided foundational insights into Temporal IR. Alonso et al. (2011) discusses challenges such as real-time streams, exploratory temporal search, and spatio-temporal retrieval. Campos et al. (2014) offers a broad overview of document dating, time-aware ranking, and query understanding, covering both explicit and implicit time signals. Kanhabua and Anand (2016) complements these with a tutorial on temporal indexing and ranking, emphasizing the detection of temporal query intent. There has been however no recent systematic overview of the field despite much research interest.
As a parallel research line, TQA over knowledge graphs has gained considerable attention Jia et al. (2021); Saxena et al. (2021); Chen et al. (2023); Xiong et al. (2024). Our survey focuses on temporally aware IR/QA over text. We review both traditional and neural approaches to core tasks such as temporal tagging, event dating, time-aware retrieval, and temporal reasoning. To our knowledge, no prior survey brings together recent developments across these tasks in the context of text-based IR/QA. Other related topics, including temporal fact verification Barik et al. (2024) and timeline summarization Sojitra et al. (2024), are discussed only when directly relevant.
forked edges,
for tree=
grow=east,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=left,
font=,
rectangle,
rounded corners,
align=left,
minimum width=0.1em,
edge+=darkgray, line width=0.8pt,
s sep=2pt,
inner xsep=2pt,
inner ysep=2pt,
draw=hidden-draw,
ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center,
where level=1text width=6em,font=,,
where level=2text width=7em,font=,,
where level=3text width=7em,font=,,
where level=4text width=11.3em,font=,,
where level=5text width=6.8em,font=,,
Temporal IR/QA, draw=black, fill=gray!10, thick, text=black, ver
Dataset &
Benchmarks ( §3),
color=orange!90, fill=orange!20, thick, text=black
Temporal Document
Collections ( §3.1),
color=orange!90, fill=orange!20, thick, text=black
Diachronic Collection,
, color=orange!90, fill=orange!20, thick, text=black,
Sandhaus (2008), Silcock et al. (2024),
Trischler et al. (2017)
,color=orange!90, fill=orange!20, thick, text=black, leaf
Synchronous Collection,
color=orange!90, fill=orange!20, thick, text=black
Vrandečić and Krötzsch (2014)
,color=orange!90, fill=orange!20, thick, text=black, leaf
TQA Datasets ( §3.2),
color=orange!90, fill=orange!20, thick, text=black
Simple QA
, color=orange!90, fill=orange!20, thick, text=black,
ArchivalQA Wang et al. (2022),
TimeQA Chen et al. (2021),
ChroniclingAmericaQA Piryani et al. (2024b),
SituatedQA Zhang and Choi (2021)
,color=orange!90, fill=orange!20, thick, text=black, leaf
Complex QA
, color=orange!90, fill=orange!20, thick, text=black,
ComplexTempQA Gruber et al. (2024),
Complex-TR Tan et al. (2024),
TempLAMA Dhingra et al. (2022),
MenatQA Wei et al. (2023),
TempReason Tan et al. (2023),
TempTabQA Gupta et al. (2023)
, color=orange!90, fill=orange!20, thick, text=black, leaf
TIR Dataset ( §3.3),
color=orange!90, fill=orange!20, thick, text=black
Joho et al. (2014), Joho et al. (2016),
,color=orange!90, fill=orange!20, thick, text=black, leaf,text width=20em
Temporal Tasks ( §4),
color=cyan!100, fill=cyan!15, thick, text=black
Prediction,
color=cyan!100, fill=cyan!15, thick, text=black
Event Dating,
color=cyan!100, fill=cyan!15, thick, text=black,
Das et al. (2017), Wang et al. (2021b)
Morbidoni et al. (2018),
Wang et al. (2021b)
, color=cyan!100, fill=cyan!15, thick, text=black, leaf
Document Dating,
color=cyan!100, fill=cyan!15, thick, text=black,
Dalli (2006), Jatowt et al. (2007)
Kumar et al. (2012)
,, color=cyan!100, fill=cyan!15, thick, text=black, leaf
Focus Time Estimation,
color=cyan!100, fill=cyan!15, thick, text=black,
Jatowt et al. (2013),
Shrivastava et al. (2017)
,, color=cyan!100, fill=cyan!15, thick, text=black, leaf
Query Time Profiling,
color=cyan!100, fill=cyan!15, thick, text=black,
Kanhabua and Nørvåg (2010),
Dakka et al. (2008),
Jones and Diaz (2007)
, color=cyan!100, fill=cyan!15, thick, text=black, leaf
Approaches ( §5),
color=red!80, fill=red!15, thick, text=black
Rule-Based &
Statistical Models ( §5.1),
color=red!80, fill=red!15, thick, text=black
Li and Croft (2003), Alonso et al. (2007), Alonso et al. (2009), Arikan et al. (2009),
Metzler et al. (2009),
Kanhabua et al. (2012), Jatowt et al. (2005),
Holzmann and Anand (2016), Kulkarni et al. (2011) , Harabagiu and Bejan (2005),
Saquete et al. (2004), Saquete et al. (2009),
Setty et al. (2017), Anand et al. (2011), Anand et al. (2012),
color=red!80, fill=red!15, thick, text=black, leaf, text width=20em
Temporal Language
Models ( §5.2),
color=red!80, fill=red!15, thick, text=black
Dhingra et al. (2022), Rosin et al. (2022), Wang et al. (2023)
Ren et al. (2023), Cole et al. (2023), Son and Oh (2023), Cole et al. (2023),
Rosin and Radinsky (2022), Su et al. (2023),
,color=red!80, fill=red!15, thick, text=black, text width=20em, leaf
Temporal RAG ( §5.2.1),
color=red!80, fill=red!15, thick, text=black
Gade and Jetcheva (2024), Wu et al. (2024), Wang et al. (2023)
Abdallah et al. (2025), Siyue et al. (2024), Qian et al. (2024), Vu et al. (2024),
,color=red!80, fill=red!15, thick, text=black, text width=20em, leaf
Temporal Reasoning ( §5.2.2),
color=red!80, fill=red!15, thick, text=black
Han et al. (2021), Niu et al. (2024), Mathur et al. (2021)
Yang et al. (2020), Wang and Zhao (2024), Zhang et al. (2024), Feng et al. (2023)
,color=red!80, fill=red!15, thick, text=black, text width=20em, leaf
Appendix B Temporal Processing Concepts
We mention here other concepts broadly related to temporal processing.
Temporal taggers are essential tools in temporal information processing; they identify and standardize time expressions in text, such as “March 15, 2021” or “yesterday,” converting them into formats like YYYY-MM-DD and categorizing them (e.g., DATE, DURATION). Popular taggers like HeidelTime Strötgen and Gertz (2010), SUTime Chang and Manning (2012), Temponym tagger Kuzey et al. (2016b), CogCompTime Ning et al. (2018) support a range of languages and domains, forming the foundation for downstream tasks including TQA, event ordering, and timeline construction.
Additionally, Temponyms Kuzey et al. (2016a) are free-text phrases that implicitly refer to specific time periods or events but are not recognized as standard temporal expressions, for Instance, "Greek referendum" or “Clinton’s presidency”. Recognizing and resolving these expressions is essential for comprehensive temporal understanding. Other related concepts include temporal granularity (typically ranging from day to decade), temporal proximity (the temporal closeness of a document to the query’s target time, influencing ranking), and temporal distribution patterns in retrieval results. Effectively leveraging these signals is key to building time-aware systems Campos et al. (2014).
Temporal Disambiguation resolves ambiguous time references (e.g., identifying which "Tuesday" is being discussed), addressing temporal ambiguity in both queries and documents Piryani et al. (2024a). Temporal Co-reference involves identifying and linking different mentions of the same temporal entity within or across documents, such as connecting “that year” to “2020” Ning et al. (2018). Timeline Extraction automatically constructs a chronological sequence of events or facts from text, to answer questions requiring event ordering, such as constructing a historical timeline Bedi et al. (2017).
More advanced reasoning tasks include Temporal Reasoning, which infers time-related relationships, such as determining the order of events or calculating durations between them. It is crucial for answering complex questions like "What happened in Poland after World War II and before 1960?" Leeuwenberg and Moens (2019). Temporal Aggregation synthesizes information from multiple time periods to answer broad or comparative questions (e.g., “How has climate policy evolved over the last decade?”). Temporal Robustness Wallat et al. (2025) refers to the resiliency of systems to adversarial changes in time-related elements (e.g., altering a date in a query, or its position in a sentence) in the form of temporal perturbations. It is used in evaluation to assess temporal reasoning stability.
Appendix C Temporal Prediction Tasks
Temporal prediction tasks are crucial in understanding and organizing time-sensitive textual data. Despite sharing the common objective of grounding text in time, these tasks differ in focus, granularity, and application. In this section, we explore related temporal prediction tasks—document dating, document focus time estimation, temporal query profiling, and event occurrence time estimation, which provide complementary insights and support distinct applications. Each task addresses unique aspects of temporal analysis, from inferring document creation times to profiling query intent. Below, we review these tasks, their methodologies, and key contributions, emphasizing their roles in temporal IR and QA.
C.1 Document Dating
Document dating refers to the task of estimating a document’s creation time (e.g., publication date) based on its textual content, especially when metadata is missing, unreliable, or unavailable. The input is the full document text, and the output is a timestamp, typically at year or month granularity.
Early approaches, such as that by de Jong et al. (2005), leveraged unigram language models trained over distinct time periods to determine when a document’s vocabulary was most prevalent. Building on this, Kanhabua and Nørvåg (2008) integrated additional linguistic features such as part-of-speech tags, tf-idf scores, and collocations to better capture temporal patterns. Dalli (2006) introduced an unsupervised method for automatic document dating using periodic word usage. Kumar et al. (2012) trained language models over discretized time intervals (chronons) using Wikipedia biographies. Niculae et al. (2014) model document dating as a pairwise ranking problem using logistic regression. More recently, Vashishth et al. (2018) introduced a neural method employing Graph Convolutional Networks (GCNs) to model syntactic and temporal relations jointly.
Document dating is crucial in temporal indexing, digital preservation, and metadata recovery, particularly for historical or noisy corpora. Beyond textual content analysis, several methods estimate the creation date of web resources. Jatowt et al. (2007) was the first approach for dating content of web pages. The authors estimated timestamps of individual content elements of web pages using their archived snapshots. SalahEldeen and Nelson (2013) developed Carbon Date, a tool that aggregates signals from multiple online sources, such as first tweets, archive snapshots, URL shorteners, and search engine crawls, to estimate a webpage’s creation date.
C.2 Document Focus Time Estimation
Document focus time estimation aims to identify the historical time periods that a document discusses, which may differ from its actual publication date. For example, a news article published in 2021 that analyzes the 9/11 attacks would have a focus time centered around September 2001. The input to this task is the document’s full text, and the output consists of one or more temporal intervals that represent the document’s narrative temporal scope. Jatowt et al. (2013) proposed a graph-based method that models co-occurrences between terms and dates to identify salient temporal associations within the text. Building on this, Jatowt et al. (2015) introduced a method that estimates focus time using statistical evidence from external corpora, even when explicit temporal expressions are limited. Shrivastava et al. (2017) further advanced this line of work by linking documents to Wikipedia concepts, leveraging their temporal relations to estimate focus times. This task supports historical analysis, event-centric retrieval, and timeline generation, providing insights into the temporal context of textual content.
C.3 Temporal Query Profiling
Temporal query profiling determines a query’s temporal intent and time of interest, such as whether it refers to the past, future, or is atemporal. The input is a short keyword query (e.g., "Ukraine-Russia war"), and the output is an inferred time or temporal distribution. Kanhabua and Nørvåg (2010) estimated query time by analyzing timestamps of top-k retrieved documents, while Dakka et al. (2008) and Jones and Diaz (2007) modeled temporal distributions of relevant documents. Kanhabua and Nørvåg (2011) conducted a comparative evaluation of five temporal ranking approaches (LMT, LMTU, TS, TSU, FuzzySet), evaluating their ability to model uncertainty and adapt to temporal variance. Gupta and Berberich (2014) combined timestamp metadata with temporal expressions in document content to infer precise time intervals. Temporal query profiling is essential for time-aware IR, as it enables query disambiguation, improves temporal relevance ranking, and supports applications such as event-centric search and timeline construction.
C.4 Event Occurrence Time Estimation
Event occurrence time estimation aims to predict the specific date on which an event occurred, given a short textual description (e.g., "Plane crash in Armenia kills 36"). Unlike document-centric tasks, this focuses on the event mention itself and typically requires high-granularity outputs—such as day- or month-level timestamps.
Das et al. (2017) introduced time vectors combining word and global temporal embeddings, estimating dates via cosine similarity. Morbidoni et al. (2018) leveraged structured knowledge bases such as DBpedia and Wikipedia to link event descriptions to temporally grounded entities. Honovich et al. (2020) proposed a neural approach with sentence extraction, LSTM with attention, and an MLP classifier for date prediction. More recently, Wang et al. (2021b) introduced TEP-Trans, a Transformer-based model that formulates event time prediction as a multivariate time series forecasting problem using features extracted from temporal news collections.
Summary: While these temporal prediction tasks are highly interrelated, each aiming to anchor textual information within a temporal context, they address distinct facets of temporal understanding. Document dating predicts when a document was created, whereas document focus time estimation identifies when the content is about, which may precede or differ from the creation time. Temporal query profiling focuses on the user’s intent, inferring when the query is directed in time rather than analyzing any specific document. Finally, event occurrence time estimation deals with precise, often fine-grained dating of event mentions, requiring models to infer real-world event timelines from sparse input. Together, these tasks form a complementary suite of temporal reasoning capabilities, enabling robust time-aware information retrieval and question answering systems.