### v3

# It’s High Time : A Survey of Temporal Question Answering

Abstract
Time plays a critical role in how information is generated, retrieved, and interpreted. In this survey, we provide a comprehensive overview of Temporal Question Answering (TQA), a research area that focuses on answering questions involving temporal constraints or context. As the amount of time-stamped content from sources like news articles, web archives, and knowledge bases increases, systems must address challenges such as detecting temporal intent, normalizing time expressions, ordering events, and reasoning over evolving or ambiguous facts. We focus on recent advances in TQA enabled by neural architectures, especially transformer-based models and Large Language Models (LLMs), highlighting progress in temporal language modeling, retrieval-augmented generation (RAG), and temporal reasoning. We also discuss benchmark datasets and evaluation strategies designed to test temporal robustness, recency awareness, and generalization.
It’s High Time
: A Survey of Temporal Question Answering
Bhawna Piryani,♠ Abdelrahman Abdallah,♠ Jamshid Mozafari,♠ Avishek Anand,♡ Adam Jatowt♠
♠University of Innsbruck
{bhawna.piryani, abdelrahman.abdallah,
jamshid.mozafari, adam.jatowt}@uibk.ac.at
♡TU Delft
avishek.anand@tudelft.nl
1 Introduction
From analyzing centuries-old texts to answering questions about emerging developments, time shapes how we seek and interpret information. As digital content grows across time-stamped sources like news archives, social media, and knowledge bases, the ability to process and reason over temporal information has become essential Alonso et al. (2007). Temporal Question Answering (TQA) is a subfield of QA that focuses on answering questions constrained by or dependent on time, requiring systems to go beyond surface-level retrieval and processing and engage in temporal understanding and reasoning.
TQA faces distinct challenges that set it apart from standard QA tasks. These include temporal ambiguity resolution, where vague expressions like "recently" or "after the war" must be interpreted relative to the document or query context; cross-temporal reasoning, which requires understanding causal and sequential relationships between events across time; and knowledge volatility, where the continual evolution of facts renders static corpora and pre-trained models inadequate for up-to-date temporal reasoning. Temporal intent may be expressed explicitly (e.g., "post-World War II", "in 1998") or implicitly, requiring systems to infer appropriate timeframes from context. Effectively addressing these challenges demands advanced temporal reasoning, context-aware retrieval strategies, and mechanisms for dynamically adapting to changing knowledge (Berberich et al., 2010).
These challenges are illustrated in Figure 1, which presents two example questions. Q1: "At what age did Obama win the Nobel Peace Prize?" requires identifying and grounding two temporal anchors, Obama’s birth year (1961) and the year he received the Nobel Peace Prize (2009), followed by computing the interval between these two events to derive the answer: 48 years old.
Q2: "What does President Obama’s climate policy tell us about how the U.S. viewed climate change during his late years of service?" requires a more nuanced understanding of temporal context. The phrase "late years of service" requires anchoring to Obama’s presidency timeline (2009-2017). Accurate answering necessitates accessing contemporaneous policy documents from that specific period, rather than relying on retrospective analyses, which highlights how the temporal context determines which knowledge sources are relevant. As shown in the figure, the model must interpret relative expressions such as "today" or "next week" that appear in document content, anchoring them to the document’s publication date (August 4, 2015) to properly contextualize Obama’s climate policy statements within their historical moment.
Research in TQA has evolved significantly, transitioning from rule-based pipelines Harabagiu and Bejan (2005) and statistical models Berberich et al. (2010); Wang et al. (2020) to modern neural systems capable of complex temporal reasoning Dhingra et al. (2022); Wang et al. (2023). Early approaches relied on handcrafted rules to normalize time expressions and resolve event sequences, but lacked scalability and struggled with diverse temporal contexts. The emergence of pre-trained language models has transformed the field by enabling robust temporal reasoning Jain et al. (2023), temporal event sequencing Lin et al. (2021), and adaptation to evolving knowledge through continual learning Han et al. (2021). These advancements paved the way for more dynamic systems that perform multi-hop temporal inference and incorporate up-to-date information through retrieval augmentation.
Despite this progress, current systems still face challenges such as handling future-oriented questions, reasoning over temporally inconsistent documents, and mitigating knowledge decay. Addressing these problems requires a deeper understanding of the field’s current capabilities and limitations.
While prior surveys111See Section A for a detailed discussion. have addressed general QA/IR techniques Kolomiyets and Moens (2011); Zhu et al. (2023, 2021) or narrow aspects of temporal processing Kobayashi and Takeda (2000); Kanhabua et al. (2015); Campos et al. (2014), a dedicated and up-to-date synthesis of TQA is missing. The most recent work to partially cover TQA appeared nearly a decade ago Campos et al. (2014), before the rise of pre-trained language models and large-scale temporal benchmarks that now define the field.
In this survey, we present a comprehensive synthesis of recent work in TQA from an unstructured text. Our scope is limited to models and datasets that operate over natural language questions and unstructured documents, excluding structured approaches based on knowledge graphs (KGs); for KG-based methods, readers may refer to Su et al. (2024a); Saxena et al. (2021). Similarly, while temporal reasoning is a component of TQA, comprehensive coverage of temporal reasoning datasets and benchmarks is beyond our scope; for extensive analysis of temporal reasoning capabilities, we refer readers to Wenzel and Jatowt (2023).
We review the evolution of TQA methods, benchmark datasets, evaluation strategies, and emerging challenges. To guide this discussion, we present a taxonomy, as illustrated in Figure 2 in the Appendix, categorizing TQA tasks, datasets, and approaches. Our goal is to provide a unified framework that synthesizes recent progress, clarifies open questions, and supports the design of next-generation time-aware systems.
2 Key Concepts
We first introduce the core concepts related to TQA. Temporal Information Retrieval (TIR) aims to retrieve documents that are not only topically relevant but also aligned with the query’s temporal intent. This intent may be explicit, such as “Olympics 2024”, or implicit, such as “latest Apple earnings”. TIR relies on different temporal signals, including document timestamps (publication dates), temporal expressions (“March 2023”), and event mentions (“2024 Olympics”) to assess a document’s temporal relevance, which indicates how well its content aligns with the query’s intended time frame Kanhabua and Nørvåg (2008); Singh et al. (2016). While TIR facilitates time-sensitive access to documents, it often lacks mechanisms for performing deeper temporal reasoning over the content itself.
On the other hand, Temporal Question Answering (TQA) focuses on answering questions with temporal constraints, either explicitly stated, such as "Who won the Nobel Prize in Physics in 2020?", or implied, for instance, "What are the latest US climate policies?". Success in TQA requires understanding the question’s temporal intent and retrieving documents relevant to the corresponding time frame, or published around that time. Unlike TIR, TQA often involves more sophisticated processing, including interpreting temporal expressions, ordering events, and grounding answers in the appropriate time frame. As a result, it frequently requires multi-hop reasoning across documents and temporal signals.
TQA relies on diverse temporal cues. Temporal signals include explicit temporal expressions like "March 2023", relative expressions such as "last week", implicit cues like "recently", and event-based references like "2024 Olympics". These require contextual interpretation and sometimes external temporal knowledge. Temporal metadata, such as document timestamps, indicate publication time and often serve as proxies for judging the freshness of content.
A critical yet often underutilized concept is that of document focus time, the time period a document primarily discusses, which may differ from its publication timestamp. For example, a news article published in 2013 about the 2010 Academy Awards has a publication date of 2013 but a focus time of 2010. This distinction is crucial in TQA, where users often seek information tied to specific historical or future contexts rather than content merely anchored to when it was written.
Estimating focus time is non-trivial, often requiring temporal expression normalization, burst detection, or event timestamping. When accurately modeled, it improves retrieval quality in time-sensitive settings, especially for long-range archives or temporally anchored questions. Prior work has shown that focus time estimation can enhance performance in tasks such as historical QA, event ordering, and temporally grounded search Jatowt et al. (2013); Wang et al. (2020).
In Appendix B we discuss other related concepts, such as temporal taggers, temponyms, granularity, temporal reasoning, timeline extraction, disambiguation, and robustness.
Dataset #Questions Knowledge Source Creation Method Answer Type Time Frame Temporal Metadata Multi-Hop NewsQA Trischler et al. (2017) 119k News CS Freeform 2007-2015 ✗ ✗ TDDiscourse Naik et al. (2019) 6.1k News CS Extractive Unspecified ✗ ✗ TORQUE Ning et al. (2020) 21k News CS Abstractive - ✗ ✗ ArchivalQA Wang et al. (2022) 532k News AG Extractive 1987-2007 ✓ ✗ TimeQA Chen et al. (2021a) 41.2K Wikipedia AG Extractive 1367-2018 ✗ ✗ TiQ Jia et al. (2024) 10K Wikipedia AG Freebase Unspecified ✗ ✗ TempQuestions Jia et al. (2018) 1.2k Freebase AG Extractive Unspecified ✗ ✓ TemporalQuestions Wang et al. (2021a) 1K News CS Extrcative 1987-2007 ✓ ✗ TempLAMA Dhingra et al. (2022) 50k News CS Extractive 2010-2020 ✓ ✗ ComplexTempQA Gruber et al. (2024) 100,228k Wikipedia AG Extractive 1987-2023 ✓ ✓ MenatQA Wei et al. (2023) 2.8k Wikipedia AG Extractive 1367-2018 ✗ ✗ PAT-Question Meem et al. (2024) 6,1k Wikipedia CS Extractive - ✗ ✓ TempTabQA Gupta et al. (2023) 11.4k Wikipedia Info box CS Abstractive - ✗ ✗ SituatedQA Zhang and Choi (2021) 12.2k Wikipedia CS - 2021 ✗ ✗ UnSeenTimeQA Uddin et al. (2024) 3.6k Synthetic AG Abstractive - ✗ ✓ ChroniclingAmericaQA Piryani et al. (2024b) 485k News AG Extractive 1800-1920 ✓ ✗ FRESHQA Vu et al. (2024) 600 Google Search CS - - ✗ ✓ COTEMPQA Su et al. (2024b) 4.7k Wikidata CS Abstractive 2023 ✗ ✓ Test of Time (ToT) Fatemi et al. (2024) 1.8k Synthetic AG Abstractive - ✗ ✓ TIMEDAIL Qin et al. (2021) 1.1k DailyDialog CS Multiple-choice - ✗ ✗ Complex-TR Tan et al. (2024) 10.8 Wikipedia+Google Search AG Multi-answer 2023 ✗ ✓ StreamingQA Liska et al. (2022) 147k News Cs Extractive 2007-2020 ✓ ✓ TRACIE Zhou et al. (2021) 5.4k Wikipedia CS abstractive 2020 ✗ ✗ ForecastQA Jin et al. (2021) 10.3k News CS Multiple-Choice 2015-2019 ✓ ✓ TEMPREASON Tan et al. (2023) 52.8k Wikipedia/Wikidata SC Abstractive 634-2023 ✗ ✗ TemporalAlignmentQA Zhao et al. (2024) 20k Wikipedia AG Abstractive 2000-2023 ✗ ✗ ReaLTimeQA Kasai et al. (2023) 5.1k Search CS Multiple-choice 2020-2024 ✗ ✗
3 Datasets and Evaluation Benchmarks
Building on the temporal representation concepts introduced in Section 2, we now examine the resources and benchmarks used to evaluate systems that address temporally grounded questions. The development of TQA systems fundamentally depends on the availability of temporally grounded datasets and robust evaluation methodologies used for training, testing, and benchmarking time-aware models. This section provides an overview of temporal datasets, organized into two categories: temporal document collections that serve as knowledge sources, and TQA datasets that enable systematic evaluation of temporal reasoning capabilities.
3.1 Temporal Document Collections
Prior work has utilized three types of temporally structured document collections, each offering distinct advantages for temporal reasoning tasks.
Diachronic corpora consist of time-stamped documents spanning extensive time periods. They support retrospective retrieval, diachronic analysis, and event-based reasoning. Prominent examples include the New York Times Annotated Corpus (1987–2007; 1.8m articles) Sandhaus (2008), which, for example, serves as the basis for ArchivalQA Wang et al. (2022) dataset, and the CNN/Daily Mail corpus (2007–2015; 313k articles) Hermann et al. (2015) used, e.g., in NewsQA Trischler et al. (2017). The Chronicling America collection (1800–1920) offers digitized historical newspaper articles and supports long-range historical QA via ChroniclingAmericaQA Piryani et al. (2024b). More recently, the Newswire corpus Silcock et al. (2024) has expanded the length of time frames, providing 2.7 million newswire articles published between 1878 and 1977. It is enriched with metadata including geo-referenced datelines, Wikipedia/Wikidata entity links, and topical annotations, enabling fine-grained historical and spatio-temporal modeling. Another widely used corpus is CUSTOMNEWS Lazaridou et al. (2021) (1969–2019), which consists of crawled English news sources and spans diverse domains including politics, finance, and sports.
Diachronic corpora are also used in a range of related temporal tasks, including semantic drift detection Hamilton et al. (2016), event burst modeling Radinsky and Horvitz (2013), and timeline construction Gutehrlé et al. (2022).
Synchronic corpora represent a coherent snapshot of the world at a specific point in time. Unlike diachronic corpora, which typically span decades or years, synchronous collections capture a temporally aligned view, sometimes in conjunction with structured knowledge bases. Wikipedia articles Vrandečić and Krötzsch (2014), for example, reflect a particular version of world knowledge at a certain time (when the dump was made) and can be linked to Wikidata timestamps. Datasets like TimeQA Chen et al. (2021a), TEMPREASON Tan et al. (2023), and ComplexTempQA Gruber et al. (2024) build on Wikipedia snapshots to support temporally-scoped QA grounded in a time-specific context.
Finally, Annotated temporal corpora with explicit temporal annotations facilitate more structured forms of temporal reasoning. TimeBank Pustejovsky et al. (2003) introduced TimeML to annotate temporal expressions, events, and their temporal relations. Follow-up datasets like WikiWars Mazur and Dale (2010) and RED O’Gorman et al. (2016) extended it to historical narratives and causal relations, respectively.
These corpora serve as gold-standard resources for temporal tagging, relation extraction, and developing models that can explicitly reason about temporal structures in text.
3.2 TQA Datasets
TQA datasets enable the evaluation of how well systems can answer questions that require temporal understanding and reasoning. They vary along multiple dimensions, including Knowledge Source, Temporal Orientation, Temporal Explicitness, and Reasoning Complexity. Table 1 provides a comprehensive comparison of major TQA datasets across these dimensions.
Knowledge Source: TQA datasets are commonly derived from diachronic or synchronic corpora. Primary Source Dataset derive from diachronic corpora, providing contemporaneous accounts written when events occurred. These datasets test models’ ability to retrieve and reason over temporally anchored document collections with authentic historical perspectives. Datasets such as NewsQA Trischler et al. (2017), TDDiscourse Naik et al. (2019), TORQUE Ning et al. (2020), ArchivalQA Wang et al. (2022), TKGQA Ong et al. (2023), ChroniclingAmericaQA Piryani et al. (2024b), are curated from historical news sources and support temporal reasoning over period-specific content.
In contrast, Synchronic Corpora like Wikipedia basically constitute Secondary Sources since they provide retrospective view of the past. They have been used to build datasets like TimeQA Chen et al. (2021a), TEMPREASON Tan et al. (2023), TiQ Jia et al. (2024), and ComplexTempQA Gruber et al. (2024), which support fine-grained reasoning across temporally scoped, consistent knowledge bases.
Temporal Orientation: While most datasets focus on past events, future-oriented QA datasets remain relatively rare. Still, they are increasingly important for evaluating models’ ability to perform predictive and hypothetical reasoning. ForecastQA Jin et al. (2021), FutureContext Mutschlechner and Jatowt (2025) and TimeBench Chu et al. (2024) are among the few benchmarks that include questions about future events, testing models’ ability to perform timeline projections and forecast-based inference.
Question Type: Temporal questions can be broadly classified by their explicitness in referencing time. Datasets like TimeQA Chen et al. (2021a), SituatedQA Zhang and Choi (2021) and TempQuestions Jia et al. (2018) contain Explicit Temporal Questions with clear temporal markers, such as "What happened in 1947?", signaling temporal intent directly.
In contrast, Implicit Temporal Questions omit direct time references but still require temporal inference. For instance, "Who was Prime Minister of the UK when the Berlin Wall fell?" requires inferring the date of the event and then linking it to a temporally relevant fact. Datasets such as TiQ Jia et al. (2024) and TORQUE Ning et al. (2020) focus on implicit reasoning, testing event-event and event-time relationships. Others like ArchivalQA Wang et al. (2022), TemporalQuestions Wang et al. (2021a), and ComplexTempQA Gruber et al. (2024) combine both question types, offering a spectrum of temporal reasoning demands from explicit, time-anchored queries to implicit, event-based inference.
Temporal Reasoning Complexity: TQA tasks also vary in the depth of reasoning they require. Simple Temporal Questions typically involve direct lookups, such as identifying the date of a specific event or the state of the world at a given time. Early datasets like NewsQA Trischler et al. (2017) and TempLAMA Dhingra et al. (2022) largely belong to this category. In contrast, Complex Temporal Questions demand more intricate processing such as multi-hop reasoning, temporal filtering, or synthesizing information across events. For example, the question “What major international agreements were signed after World War I but before World War II?” necessitates multi-hop temporal reasoning and contextual comparison. Datasets like MenatQA Wei et al. (2023), TempReason Tan et al. (2023), Complex-TR Tan et al. (2024), and ComplexTempQA Gruber et al. (2024) are explicitly designed to evaluate these advanced reasoning capabilities. Others like TimeBench Chu et al. (2024) span both simple and complex reasoning levels, including tasks such as timeline construction or event duration inference.
Summary: TQA poses challenges that go beyond annotation quality or dataset scale. Unlike static QA tasks, the answers to temporal questions can shift over time, necessitating benchmarks that can account for answer drift and temporal volatility. Most existing datasets are static snapshots; notable exceptions include RealTimeQA Kasai et al. (2023) and FreshQA Vu et al. (2024), which incorporate periodic updates, though these updates are often resource-intensive. Alternatively, datasets like PATQA Meem et al. (2024) explore scalable, automated update mechanisms. A persistent issue is temporal ambiguity Piryani et al. (2024a), where missing or implicit time references hinder both annotation and evaluation. Structural biases also emerge: diachronic corpora tend to have more data surrounding major events and contain detailed information on past events, yet they reflect contemporary perspectives. In contrast, synchronic corpora provide broader coverage but often lack fine-grained event granularity. Annotation strategies further involve trade-offs: crowdsourced datasets are typically small but high-quality, whereas automatically generated datasets are larger but noisier. Moreover, most benchmarks fail to isolate specific reasoning skills such as duration inference or event ordering hereby limiting diagnostic evaluation.
While this section focuses on general-domain datasets, specialized domains such as medical, legal, and financial contexts present unique temporal reasoning challenges and dataset requirements, as shown in Appendix E. To support meaningful progress, robust TQA datasets must evolve over time, resolve temporal ambiguity, and facilitate fine-grained reasoning assessment.
4 Approaches in Temporal QA
A wide range of approaches have been developed to address the challenges of TQA, from early rule-based systems and statistical models to neural networks and large language models (LLMs). Tables 2 and 3 (see Appendix) provide a comprehensive comparison of these temporal QA approaches across architectural paradigms, temporal representations, and methodological strategies. They differ in how they represent temporal information, reason over temporal relationships, and adapt to changing world knowledge.
Effective TQA often relies on several core Temporal Prediction Tasks that support deeper temporal understanding. These tasks focus on inferring implicit or missing temporal signals from text, which is especially important when explicit temporal metadata is sparse or unreliable. Key tasks include Event Dating, Document Dating, Focus Time Estimation, Query Time Profiling, and Event Occurrence Prediction. They enable a range of downstream applications such as timeline construction, historical information retrieval, and time-sensitive question answering.
In this section, we focus on neural approaches, particularly temporal language models and retrieval-augmented generation (RAG) for TQA. For a more detailed discussion of temporal prediction tasks, as well as traditional rule-based and statistical QA methods, we refer the reader to Appendix C.
4.1 Temporal Language Models
The emergence of deep learning has significantly advanced TQA by enabling models to capture temporal dependencies and contextual nuances. Recent research has led to the development of Temporal Language Models (TLMs) that explicitly incorporate temporal signals during pretraining or fine-tuning. Models such as TempoT5 Dhingra et al. (2022), TempoBERT Rosin et al. (2022), and BiTimeBERT Wang et al. (2023) included timestamps and temporal expressions directly into their training inputs or used time-focused pretraining tasks. These strategies improve temporal generalization in downstream tasks such as semantic change detection and TQA.
Beyond timestamp inputs, more recent work introduces architectural modifications to capture temporal structure internally. TALM Ren et al. (2023) models time hierarchically, learning both short-term and long-term temporal word representations. SG-TLM Su et al. (2023) enhances temporal sensitivity by masking syntactic spans associated with time expressions, encouraging the model to learn implicit temporal cues. Similarly, models like Temporal Span Masking (TSM) Cole et al. (2023) and Temporal Attention Rosin and Radinsky (2022) explicitly incorporate temporal supervision into attention heads, enabling more fine-grained alignment between time expressions and relevant content.
Some TLMs focus on generation tasks that require temporal grounding. For instance, Cao and Wang (2022) proposes temporal prompts, both textual and continuous vector-based, that guide generation with time-specific context. TCQA Son and Oh (2023) introduces a synthetic QA dataset and a span-selection task that aligns answers with their temporal context, simulating historical reasoning and timeline consistency.
A common trend across TLMs is the shift from treating time as an auxiliary signal to making it a core component of the modeling process. Timestamp conditioning, temporal pretraining tasks and attention adaptation have improved temporal reasoning, but most models still rely on static corpora and struggle with vague or implicit expressions. TLMs represent a significant step forward but require further development to handle real-world complexity and temporal drift.
4.2 Temporal RAG
While TLMs improve temporal understanding through pretraining, they remain limited by the static nature of training data. To address evolving information needs and reduce temporal hallucinations, recent work has turned to Retrieval-Augmented Generation (RAG) that integrates neural retrieval with generation to incorporate up-to-date, time-relevant evidence at inference time.
A key trend in this space is the integration of temporal signals directly into the retrieval process. Models such as TempRetriever Abdallah et al. (2025) and TsContriever Wu et al. (2024) extend dense retrievers to account for temporal relevance by encoding queries and passages with timestamp-aware embeddings. These approaches improve alignment between temporally scoped questions and contextually valid evidence. Similarly, TempRALM Gade and Jetcheva (2024) enhances dense retrieval with temporal constraints to reduce factual drift and improve recency-grounded responses.
Other systems modify the retrieval generation pipeline itself. TimeR4 Qian et al. (2024) proposes a four-stage Retrieve-Rewrite-Retrieve-Rerank architecture that transforms underspecified temporal queries into explicitly time-anchored formulations before retrieving and re-ranking documents based on temporal fit. MRAG Siyue et al. (2024) addresses multi-hop reasoning across events by retrieving from multiple time-scoped sources and aggregating cross-temporal evidence for complexQA.
These models illustrate an emerging paradigm shift: from static temporal representation to real-time, adaptive temporal reasoning. Temporal RAG systems help mitigate hallucinations and better reflect recent developments, addressing temporal incompleteness. However, their success is tied to the quality and temporal alignment of the retriever, and they raise new challenges around evidence reliability and temporal grounding.
4.3 Temporal Reasoning Capabilities
While Temporal Language Models improve time-aware representation and retrieval, many TQA tasks require explicit temporal reasoning, such as tracking event sequences, resolving durations, handling temporal arithmetic, and drawing logical inferences across time. These capabilities are essential for questions involving change over time, causality, or evolving narratives, where retrieval alone is insufficient. .
Recent efforts have aimed to augment the reasoning capacity of pretrained language models (PLMs) through both architectural design and specialized training objectives. For example, ECONET Han et al. (2021) focuses on continual adaptation to changing temporal contexts, preserving consistency across updates. Models like TIMERS Mathur et al. (2021) and ConTempo Niu et al. (2024) adopt structured reasoning layers for multi-hop and document-level inference, often integrating temporal graphs or symbolic abstractions to guide event understanding.
To evaluate these capabilities more rigorously, new benchmarks have emerged. TRAM Wang and Zhao (2024) assesses LLMs on a spectrum of temporal tasks, including event frequency, duration estimation, and timeline ordering. Even state-of-the-art models like GPT-4 underperform on these tasks compared to human baselines, highlighting persistent gaps. Synthetic evaluations such as Test of Time (ToT) Fatemi et al. (2024) isolate core temporal reasoning skills like logic, arithmetic, and counterfactuals, free from the confounding effects of memorization. TODAY Feng et al. (2023) and Narrative-of-Thought Zhang et al. (2024) further probe a model’s ability to reason over temporal shifts and structured narratives.
Beyond accuracy, robustness remains a critical challenge. Studies such as Wallat et al. (2024, 2025) expose temporal blind spots in LLMs when applied on adversarial datasets, showing how small changes in timestamps or event order can significantly degrade performance, revealing a lack of temporal generalization and an over-reliance on surface form cues.
While temporal reasoning in PLMs has improved through targeted objectives and structured modeling, current systems still fall short in handling implicit cues, abstract reasoning, and robustness across time. Future work must explore deeper symbolic integration, stronger generalization under temporal drift, and evaluation setups that disentangle reasoning from retrieval and memorization.
Summary: TQA has progressed from rule-based pipelines to neural architectures that integrate temporal information through input representations, model internals, and dynamic retrieval.
TLMs improved contextualization by embedding timestamps and temporal cues, yet they remain constrained by static training corpora and a focus on explicit time expressions often falling short on vague, implicit, or underspecified temporal signals. Temporal RAG systems address some of these limitations by retrieving external, time-sensitive evidence at inference time, but their effectiveness depends heavily on the quality, relevance, and freshness of the retrieved content.
Recent reasoning-focused architectures go a step further, enabling event ordering and multi-hop inference, yet even state-of-the-art models exhibit temporal blind spots and struggle with robust generalization across temporal contexts.
Collectively, while each paradigm addresses specific limitations of its predecessor, none fully resolves the challenges of temporal ambiguity, drift, and reasoning. Future work must systematically target these failure modes, particularly the balance between memorized and retrieved knowledge, the handling of underspecified timeframes, and evaluation methods that isolate genuine temporal reasoning from pattern matching. Bridging these gaps is essential for building truly robust and temporally aware QA systems.
5 Future Research Directions
Despite significant advances in neural temporal reasoning, current TQA systems face fundamental limitations that require targeted research efforts. We identify the following critical areas where future work is needed to build temporally aware, trustworthy QA systems. Detailed technical challenges and potential solutions for each direction are provided in Appendix F.
Dynamic Temporal Knowledge Management.
Current TQA systems rely on static corpora, making them unable to respond to fast-evolving information needs. The temporal propagation problem, where small updates disrupt temporal dependencies across related events, exposes fundamental architectural limitations. Future systems must move beyond isolated fact updates toward scalable frameworks that track, edit, and reason over temporal dependencies in real time.
Temporally-Aware LLM Agents.
While LLMs demonstrate impressive general reasoning capabilities, they consistently underperform in temporal reasoning tasks. LLMs exhibit temporal hallucinations and fail to resolve context-dependent expressions such as "last Tuesday" or "since our previous discussion." Incorporating timeline tracking, event memory, and temporal reference resolution is essential for temporally coherent dialogue agents.
Diachronic and Synchronic Knowledge Integration.
Temporal questions often require combining time-sensitive sources showing change over time with stable sources summarizing information at specific points. Most current systems treat these separately, limiting their ability to answer questions involving both historical trends and current facts. Future systems should develop temporal alignment algorithms and cross-source reasoning frameworks.
Temporal Uncertainty and Confidence Modeling.
Many historical events have unclear start and end dates, yet most TQA systems treat all dates as exact. This mismatch creates problems when systems give confident answers about uncertain information, with issues compounding in multi-step reasoning. Future systems should handle uncertain dates and provide confidence scores for temporal answers.
Multilingual and Multimodal Temporal QA.
Temporal expressions, date formats, and cultural references vary widely across languages and modalities, yet most TQA systems are developed primarily for English textual input. Current systems struggle with diverse temporal signals due to limited cultural grounding and inadequate multimodal integration. Future research should focus on developing multilingual temporal taggers and cross-modal alignment techniques.
Implicit Temporal Intent Understanding.
Many temporal questions hide their time requirements rather than stating them directly. Current systems struggle with this ambiguity because they lack mechanisms to infer intended timeframes from contextual clues or shared knowledge. As users increasingly expect models to infer nuanced, context-sensitive answers, detecting hidden temporal assumptions becomes critical for avoiding plausible but misleading outputs.
Evaluation and Benchmarking for Temporal Reasoning.
Standard metrics like accuracy and F1 fall short in capturing temporal coherence or reasoning depth. Developing temporally-aware evaluation metrics remains an important open challenge for assessing and comparing future systems more effectively.
6 Conclusion
TQA plays a critical role in retrieving and reasoning over time-sensitive information in dynamic contexts. In this survey, we focus on recent advancements in the field, particularly modern TLMs and RAG approaches, examining core challenges including temporal expression recognition, event ordering, and implicit temporal reasoning.
Despite advances, current systems struggle with answer drift, temporal uncertainty, cultural and linguistic diversity, and limited temporal granularity. Most benchmarks rely on static snapshots, hindering real-time and future-oriented reasoning. Addressing these gaps requires better temporal representations, dynamic evaluation, adaptive learning.
Limitations
This survey aims to provide a comprehensive overview of TQA. There are a few important limitations to acknowledge.
We made our best efforts to be thorough, but it is possible that some relevant works may have been missed. We conducted an extensive literature review using forward and backward snowballing techniques, with particular attention to papers published in major venues such as ACL, SIGIR, EMNLP, NeurIPS, ECIR, and preprints on arXiv. On the other hand, due to page limitations, we provide only a brief summary of each method, without providing exhaustive technical details.
Appendix A Related Surveys
Advances in temporal datasets, time-aware models, and reasoning techniques have enabled systems capable of retrieving time-relevant documents, ordering events, and answering temporally constrained questions, benefiting applications such as historical analysis, fact-checking, and intelligent assistants.
While QA have been widely surveyed, most existing reviews focus on general techniques, often neglecting temporal aspects. IR surveys emphasize ranking functions, neural retrieval models, and query understanding Li et al. (2025); Zhu et al. (2023), while QA surveys center on extractive, abstractive, or multi-hop answering over static knowledge sources Zhu et al. (2021); Mavi et al. (2024). They rarely consider temporal intent, dynamic or evolving information, or event sequencing, highlighting a key gap that remains unaddressed.
Several earlier works provided foundational insights into TIR. Alonso et al. (2011) discusses challenges such as real-time streams, exploratory temporal search, and spatio-temporal retrieval. Campos et al. (2014) offers a broad overview of document dating, time-aware ranking, and query understanding, covering both explicit and implicit time signals. Kanhabua and Anand (2016) complements these with a tutorial on temporal indexing and ranking, emphasizing the detection of temporal query intent. There has been however no recent systematic overview of the field despite much research interest.
As a parallel research line, TQA over knowledge graphs has gained considerable attention Jia et al. (2021); Saxena et al. (2021); Chen et al. (2023b); Xiong et al. (2024b). Our survey focuses on temporally aware QA over text. We review both traditional and neural approaches to core tasks such as temporal tagging, event dating, time-aware retrieval, and temporal reasoning. To our knowledge, no prior survey brings together recent developments across these tasks in the context of text-based QA. Other related topics, including temporal fact verification Barik et al. (2024) and timeline summarization Sojitra et al. (2024), are discussed only when directly relevant.
forked edges,
for tree=
grow=east,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=left,
font=,
rectangle,
rounded corners,
align=left,
minimum width=0.1em,
edge+=darkgray, line width=0.8pt,
s sep=2pt,
inner xsep=2pt,
inner ysep=2pt,
draw=hidden-draw,
ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center,
where level=1text width=6em,font=,,
where level=2text width=7em,font=,,
where level=3text width=7em,font=,,
where level=4text width=11.3em,font=,,
where level=5text width=6.8em,font=,,
Temporal QA, draw=black, fill=gray!10, thick, text=black, ver
Dataset &
Benchmarks ( §3),
color=orange!90, fill=orange!20, thick, text=black
Temporal Document
Collections ( §3.1),
color=orange!90, fill=orange!20, thick, text=black
Diachronic Collection,
, color=orange!90, fill=orange!20, thick, text=black,
Sandhaus (2008), Silcock et al. (2024),
Trischler et al. (2017)
,color=orange!90, fill=orange!20, thick, text=black, leaf
Synchronic Collection,
color=orange!90, fill=orange!20, thick, text=black
Vrandečić and Krötzsch (2014)
,color=orange!90, fill=orange!20, thick, text=black, leaf
TQA Datasets ( §3.2),
color=orange!90, fill=orange!20, thick, text=black
Simple QA
, color=orange!90, fill=orange!20, thick, text=black,
ArchivalQA Wang et al. (2022),
TimeQA Chen et al. (2021a),
ChroniclingAmericaQA Piryani et al. (2024b),
SituatedQA Zhang and Choi (2021)
,color=orange!90, fill=orange!20, thick, text=black, leaf
Complex QA
, color=orange!90, fill=orange!20, thick, text=black,
ComplexTempQA Gruber et al. (2024),
Complex-TR Tan et al. (2024),
TempLAMA Dhingra et al. (2022),
MenatQA Wei et al. (2023),
TempReason Tan et al. (2023),
TempTabQA Gupta et al. (2023)
, color=orange!90, fill=orange!20, thick, text=black, leaf
Temporal Tasks ( §C),
color=cyan!100, fill=cyan!15, thick, text=black
Event Dating,
color=cyan!100, fill=cyan!15, thick, text=black,
Das et al. (2017), Wang et al. (2021b) , Morbidoni et al. (2018), Wang et al. (2021b)
, color=cyan!100, fill=cyan!15, thick, text=black, text width=20em, leaf
Document Dating,
color=cyan!100, fill=cyan!15, thick, text=black,
Dalli (2006), Jatowt et al. (2007), Kumar et al. (2012)
, color=cyan!100, fill=cyan!15, thick, text=black, text width=20em, leaf
Focus Time Estimation,
color=cyan!100, fill=cyan!15, thick, text=black,
Jatowt et al. (2013), Shrivastava et al. (2017)
, color=cyan!100, fill=cyan!15, thick, text=black, text width=20em, leaf
Query Time Profiling,
color=cyan!100, fill=cyan!15, thick, text=black,
Kanhabua and Nørvåg (2010), Dakka et al. (2008), Jones and Diaz (2007)
, color=cyan!100, fill=cyan!15, thick, text=black, text width=20em, leaf
Approaches ( §4),
color=red!80, fill=red!15, thick, text=black
Rule-Based &
Statistical Models ( §D),
color=red!80, fill=red!15, thick, text=black
Li and Croft (2003), Alonso et al. (2007), Alonso et al. (2009), Arikan et al. (2009),
Metzler et al. (2009),
Kanhabua et al. (2012), Jatowt et al. (2005),
Holzmann and Anand (2016), Kulkarni et al. (2011) , Harabagiu and Bejan (2005),
Saquete et al. (2004), Saquete et al. (2009),
Setty et al. (2017), Anand et al. (2011), Anand et al. (2012),
color=red!80, fill=red!15, thick, text=black, leaf, text width=20em
Temporal Language
Models ( §4.1),
color=red!80, fill=red!15, thick, text=black
Dhingra et al. (2022), Rosin et al. (2022), Wang et al. (2023)
Ren et al. (2023), Cole et al. (2023), Son and Oh (2023), Cole et al. (2023),
Rosin and Radinsky (2022), Su et al. (2023),
,color=red!80, fill=red!15, thick, text=black, text width=20em, leaf
Temporal RAG ( §4.2),
color=red!80, fill=red!15, thick, text=black
Gade and Jetcheva (2024), Wu et al. (2024), Wang et al. (2023)
Abdallah et al. (2025), Siyue et al. (2024), Qian et al. (2024), Vu et al. (2024),
,color=red!80, fill=red!15, thick, text=black, text width=20em, leaf
Temporal Reasoning ( §4.3),
color=red!80, fill=red!15, thick, text=black
Han et al. (2021), Niu et al. (2024), Mathur et al. (2021)
Yang et al. (2020), Wang and Zhao (2024), Zhang et al. (2024), Feng et al. (2023)
,color=red!80, fill=red!15, thick, text=black, text width=20em, leaf
Appendix B Temporal Processing Concepts
We mention here other concepts broadly related to temporal processing.
Temporal taggers are essential tools in temporal information processing; they identify and standardize time expressions in text, such as “March 15, 2021” or “yesterday,” converting them into formats like YYYY-MM-DD and categorizing them (e.g., DATE, DURATION). Popular taggers like HeidelTime Strötgen and Gertz (2010), SUTime Chang and Manning (2012), Temponym tagger Kuzey et al. (2016b), CogCompTime Ning et al. (2018) support a range of languages and domains, forming the foundation for downstream tasks including TQA, event ordering, and timeline construction.
Additionally, Temponyms Kuzey et al. (2016a) are free-text phrases that implicitly refer to specific time periods or events but are not recognized as standard temporal expressions, for Instance, "Greek referendum" or “Clinton’s presidency”. Recognizing and resolving these expressions is essential for comprehensive temporal understanding. Other related concepts include temporal granularity (typically ranging from day to decade), temporal proximity (the temporal closeness of a document to the query’s target time, influencing ranking), and temporal distribution patterns in retrieval results. Effectively leveraging these signals is key to building time-aware systems Campos et al. (2014).
Temporal Disambiguation resolves ambiguous time references (e.g., identifying which "Tuesday" is being discussed), addressing temporal ambiguity in both queries and documents Piryani et al. (2024a). Temporal Co-reference involves identifying and linking different mentions of the same temporal entity within or across documents, such as connecting “that year” to “2020” Ning et al. (2018). Timeline Extraction automatically constructs a chronological sequence of events or facts from text, to answer questions requiring event ordering, such as constructing a historical timeline Bedi et al. (2017).
More advanced reasoning tasks include Temporal Reasoning, which infers time-related relationships, such as determining the order of events or calculating durations between them. It is crucial for answering complex questions like "What happened in Poland after World War II and before 1960?" Leeuwenberg and Moens (2019). Temporal Aggregation synthesizes information from multiple time periods to answer broad or comparative questions (e.g., “How has climate policy evolved over the last decade?”). Temporal Robustness Wallat et al. (2025) refers to the resiliency of systems to adversarial changes in time-related elements (e.g., altering a date in a query, or its position in a sentence) in the form of temporal perturbations. It is used in evaluation to assess temporal reasoning stability.
Appendix C Temporal Prediction Tasks
Temporal prediction tasks are crucial in understanding and organizing time-sensitive textual data. Despite sharing the common objective of grounding text in time, these tasks differ in focus, granularity, and application. In this section, we explore related temporal prediction tasks—document dating, document focus time estimation, temporal query profiling, and event occurrence time estimation, which provide complementary insights and support distinct applications. Each task addresses unique aspects of temporal analysis, from inferring document creation times to profiling query intent. Below, we review these tasks, their methodologies, and key contributions, emphasizing their roles in temporal IR and QA.
C.1 Document Dating
Document dating refers to the task of estimating a document’s creation time (e.g., publication date) based on its textual content, especially when metadata is missing, unreliable, or unavailable. The input is the full document text, and the output is a timestamp, typically at year or month granularity.
Early approaches, such as that by de Jong et al. (2005), leveraged unigram language models trained over distinct time periods to determine when a document’s vocabulary was most prevalent. Building on this, Kanhabua and Nørvåg (2008) integrated additional linguistic features such as part-of-speech tags, tf-idf scores, and collocations to better capture temporal patterns. Dalli (2006) introduced an unsupervised method for automatic document dating using periodic word usage. Kumar et al. (2012) trained language models over discretized time intervals (chronons) using Wikipedia biographies. Niculae et al. (2014) model document dating as a pairwise ranking problem using logistic regression. More recently, Vashishth et al. (2018) introduced a neural method employing Graph Convolutional Networks (GCNs) to model syntactic and temporal relations jointly.
Document dating is crucial in temporal indexing, digital preservation, and metadata recovery, particularly for historical or noisy corpora. Beyond textual content analysis, several methods estimate the creation date of web resources. Jatowt et al. (2007) was the first approach for dating content of web pages. The authors estimated timestamps of individual content elements of web pages using their archived snapshots. SalahEldeen and Nelson (2013) developed Carbon Date, a tool that aggregates signals from multiple online sources, such as first tweets, archive snapshots, URL shorteners, and search engine crawls, to estimate a webpage’s creation date.
C.2 Document Focus Time Estimation
Document focus time estimation aims to identify the historical time periods that a document discusses, which may differ from its actual publication date. For example, a news article published in 2021 that analyzes the 9/11 attacks would have a focus time centered around September 2001. The input to this task is the document’s full text, and the output consists of one or more temporal intervals that represent the document’s narrative temporal scope. Jatowt et al. (2013) proposed a graph-based method that models co-occurrences between terms and dates to identify salient temporal associations within the text. Building on this, Jatowt et al. (2015) introduced a method that estimates focus time using statistical evidence from external corpora, even when explicit temporal expressions are limited. Shrivastava et al. (2017) further advanced this line of work by linking documents to Wikipedia concepts, leveraging their temporal relations to estimate focus times. This task supports historical analysis, event-centric retrieval, and timeline generation, providing insights into the temporal context of textual content.
C.3 Temporal Query Profiling
Temporal query profiling determines a query’s temporal intent and time of interest, such as whether it refers to the past, future, or is atemporal. The input is a short keyword query (e.g., "Ukraine-Russia war"), and the output is an inferred time or temporal distribution. Kanhabua and Nørvåg (2010) estimated query time by analyzing timestamps of top-k retrieved documents, while Dakka et al. (2008) and Jones and Diaz (2007) modeled temporal distributions of relevant documents. Kanhabua and Nørvåg (2011) conducted a comparative evaluation of five temporal ranking approaches (LMT, LMTU, TS, TSU, FuzzySet), evaluating their ability to model uncertainty and adapt to temporal variance. Gupta and Berberich (2014) combined timestamp metadata with temporal expressions in document content to infer precise time intervals. Temporal query profiling is essential for time-aware IR, as it enables query disambiguation, improves temporal relevance ranking, and supports applications such as event-centric search and timeline construction.
C.4 Event Occurrence Time Estimation
Event occurrence time estimation aims to predict the specific date on which an event occurred, given a short textual description (e.g., "Plane crash in Armenia kills 36"). Unlike document-centric tasks, this focuses on the event mention itself and typically requires high-granularity outputs—such as day- or month-level timestamps.
Das et al. (2017) introduced time vectors combining word and global temporal embeddings, estimating dates via cosine similarity. Morbidoni et al. (2018) leveraged structured knowledge bases such as DBpedia and Wikipedia to link event descriptions to temporally grounded entities. Honovich et al. (2020) proposed a neural approach with sentence extraction, LSTM with attention, and an MLP classifier for date prediction. More recently, Wang et al. (2021b) introduced TEP-Trans, a Transformer-based model that formulates event time prediction as a multivariate time series forecasting problem using features extracted from temporal news collections.
Summary: While these temporal prediction tasks are highly interrelated, each aiming to anchor textual information within a temporal context, they address distinct facets of temporal understanding. Document dating predicts when a document was created, whereas document focus time estimation identifies when the content is about, which may precede or differ from the creation time. Temporal query profiling focuses on the user’s intent, inferring when the query is directed in time rather than analyzing any specific document. Finally, event occurrence time estimation deals with precise, often fine-grained dating of event mentions, requiring models to infer real-world event timelines from sparse input. Together, these tasks form a complementary suite of temporal reasoning capabilities, enabling robust time-aware information retrieval and question answering systems.
Appendix D Rule-based & Statistical Methods
Early work in Temporal QA was dominated by rule-based systems and statistical models that laid the groundwork for core temporal tasks such as time expression normalization, event ordering, and temporal ranking. While limited in scalability and adaptability, they introduced many foundational concepts that remain relevant today.
In TIR, rule-based systems focused on extracting and normalizing time expressions to improve retrieval for time-sensitive queries Arikan et al. (2009); Alonso et al. (2007). Models like TCluster Alonso et al. (2009) and time-based language models Li and Croft (2003) used document timestamps and decay functions to model recency, while others like Berberich et al. (2010) combined metadata and vague expressions in probabilistic ranking models. To handle implicit temporal intent, techniques such as median timestamp analysis Kanhabua and Nørvåg (2010) and query log mining Metzler et al. (2009) were introduced.
Other strategies focused on enhancing recency-aware retrieval. Jatowt et al. (2005) proposed re-ranking methods using archived web snapshots to favor fresher content, while Dong et al. (2010) incorporated real-time Twitter signals, and Setty et al. (2017) used news signals into crawling and ranking to support time-sensitive queries. Efficient indexing methods were also developed to support temporal queries over evolving corpora such as Wikipedia and web archives Anand et al. (2011, 2012); Holzmann and Anand (2016). Styskin et al. (2011) introduced a machine learning model to predict recency sensitivity, combining it with greedy diversification to balance freshness and topical relevance.
As TIR matured, researchers began modeling the temporal dynamics of both queries and documents. Kulkarni et al. (2011) analyzed how user intents evolve over time, highlighting the need for adaptive retrieval strategies that can respond to temporal drift in query behavior. Joho et al. (2013) studied the prevalence of different temporal orientations of user queries, and the strategies user apply to find temporally relevant content from the past, future or present. Later systems adapted ranking strategies to temporal query profiles using machine learning Kanhabua et al. (2012) or temporal interval representations Rizzo et al. (2022).
Early QA systems like Harabagiu and Bejan (2005) relied on TimeML and lexical resources like WordNet Miller (1992) for event reasoning. To handle complex temporal questions more effectively, Saquete et al. (2004, 2009) introduced a multi-layered QA architecture that decomposed questions into temporally constrained sub-questions using temporal expression taggers like TERSEO Saquete et al. (2003). These approaches showed improved precision and generalizability across languages.
In retrospect, rule-based and statistical systems introduced many principles that remain relevant: time expression normalization, explicit modeling of document recency and user intent, and the integration of temporal structure into retrieval and reasoning pipelines. While their lack of adaptability to diverse and large-scale contexts eventually limited their impact, they provided a crucial foundation that modern neural and retrieval-augmented models continue to build upon.
Appendix E Domain-Specific Temporal QA
While much of the existing research in TQA focuses on open-domain settings using Wikipedia and news corpora, domain-specific applications present unique temporal reasoning challenges that remain largely underexplored. Unlike general TQA systems that handle diverse topics with explicit temporal markers, specialized domains require an understanding of domain-specific temporal conventions, implicit temporal relationships, and field-specific accuracy requirements.
Medical Domain
Temporal reasoning in clinical narratives has been explored in foundational work. The TimeText system, for instance, demonstrated early success in temporal clinical QA, achieving 84% accuracy on 147 time-oriented questions about discharge summaries Zhou et al. (2008). Other efforts have used semantic web techniques to answer time-based clinical questions, such as event sequencing in patient histories Tao et al. (2010).
Despite comprehensive reviews of temporal reasoning in medical text processing Olex and McInnes (2021); Zhou and Hripcsak (2007), modern temporal QA systems for medical applications remain limited. Recent work, such as Improving Health Question Answering with Reliable and Time-Aware Evidence Retrieval Vladika and Matthes (2024), incorporates recency and citation count for prioritizing reliable evidence, but falls short of full temporal reasoning. Open challenges remain in multi-hop reasoning over patient timelines, temporally scoped symptom progression, and treatment-event linking—tasks crucial to clinical decision support yet largely unexplored from a QA perspective.
Legal Domain
Legal systems present inherent temporal complexity, shaped by evolving statutes, shifting precedents, and regulatory amendments. While several legal QA datasets exist, such as Büttner and Habernal (2024); Kien et al. (2020); Chen et al. (2023a); Zhong et al. (2020), most are focused on static question answering and do not address the temporal dynamics of legal knowledge. Recent advances like ChronosLex T.y.s.s et al. (2024) introduce time-aware incremental training for better generalization across legal epochs, signaling early momentum toward temporal legal understanding.
However, most legal QA systems do not yet perform event ordering, precedent timeline alignment, or reasoning over legislative changes. For example, questions such as “Which case law was applicable prior to the 2015 amendment of the privacy statute?” remain out of scope for current systems. There is a pressing need for temporal-aware retrieval and reasoning architectures capable of navigating legal timelines and citations.
Financial Domain
In the financial domain, temporal modeling is well-established through time-series forecasting methods used for stock price prediction, volatility estimation, and risk assessment. However, these approaches typically operate on structured numerical data and do not engage in temporal reasoning over natural language. As a result, financial temporal QA remains a largely unexplored frontier.
Several financial QA datasets have emerged in recent years, primarily focused on numerical reasoning. Notable examples include FinQA Chen et al. (2021b), FinTextQA Chen et al. (2024), and FinDER Choi et al. (2025).
While these resources advance quantitative and factual QA, they fall short of supporting temporal inference. Complex questions such as “How have regulatory changes since 2008 affected current banking policies?” require systems to track policy evolution, integrate document publication dates, and synthesize trends across multiple temporal anchors (capabilities not currently addressed by existing models or datasets). Bridging this gap would require designing temporal-aware retrieval systems and multi-hop reasoning frameworks tailored to financial text, capable of answering questions grounded in historical context and regulatory shifts.
Real-World Deployment and Domain-Specific Challenges
Deploying temporal QA systems in high-stakes, real-world settings presents challenges that require specialized reasoning capabilities. In the medical domain, errors in temporal interpretation can adversely affect clinical decision-making, demanding high standards of accuracy and interpretability. Legal information systems must navigate jurisdiction-specific temporal expressions (e.g., “Pre-Brexit regulatory regime”) and shifting hierarchies of legal precedent that differ across regions and evolve over time. Financial applications require real-time temporal understanding for tasks such as regulatory compliance and market analysis, where even minor misalignments in temporal reasoning can lead to significant operational or economic consequences.
Unlike general-purpose systems that rely on clearly stated temporal cues, domain-specific QA often involves inferring timeframes from implicit, field-specific conventions such as the “perioperative period” in medicine, the “discovery phase” in legal contexts, or “earnings season” in financial reporting. These terms reflect temporally bounded phases that are well understood by domain experts but opaque to models lacking domain awareness. Addressing such challenges demands tailored temporal reasoning approaches that incorporate expert knowledge, event timelines, and contextual interpretation.
Appendix F Detailed Future Research Directions
This appendix provides a comprehensive analysis of the research challenges outlined in Section 5, detailing specific technical problems, examples, and potential research approaches for each identified direction.
F.1 Dynamic Temporal Knowledge Management
TQA systems face a persistent temporal consistency problem: their reliance on static corpora makes them unable to respond to fast-evolving information needs (Dhingra et al., 2022; Wallat et al., 2024). As real-world facts change, even small updates can disrupt temporal dependencies across related events, durations, and causal chains, a challenge known as the temporal propagation problem (Vu et al., 2024). This exposes a fundamental limitation in current architectures: they lack the modularity and temporal reasoning structure needed to adapt efficiently (Han et al., 2021). Future systems must move beyond isolated fact updates toward scalable frameworks that track, edit, and reason over temporal dependencies in real time.
For example, if a political leader’s term end’s date changes, this affects not only direct questions about their tenure but also questions about policies enacted “during their presidency” or events that occurred “before they left office.” Current systems often fail to propagate such updates, leading to inconsistencies. To address this, future systems should: (1) identify temporal dependency chains when facts change, (2) propagate updates through related facts, (3) maintain multi-hop consistency, and (4) reconcile conflicting temporal evidence. Integrated temporal knowledge graphs offer a promising direction for such dynamic reasoning.
F.2 Temporally-Aware LLM Agents
Current LLM agents face severe limitations in temporal reasoning that become pronounced in interactive settings. They consistently exhibit temporal hallucinations, generating plausible but temporally incorrect information, and fail to resolve context-dependent expressions such as "last Tuesday" or "since our previous discussion" Xiong et al. (2024a); Bazaga et al. (2025). These failures often lead to incoherent or contradictory answers across dialogue turns.
The core issue stems from architectural constraints: transformers lack persistent temporal working memory and struggle with temporal reference resolution across dialogue turns Ge et al. (2025). Unlike humans, who maintain temporal context throughout conversations, current agents treat each turn largely independently, leading to temporal inconsistencies. For instance, an agent might correctly answer a question about "last week’s events" in one turn but then fail to maintain that temporal anchor when asked a follow-up question.
Additionally, multi-turn temporal reasoning reveals deeper problems in how agents handle evolving temporal contexts. When users engage in extended conversations about temporal topics, agents often lose track of previously established temporal anchors, leading to drift in temporal understanding over the course of the dialogue.
Future agents must incorporate explicit temporal modeling mechanisms including: (1) timeline tracking systems that maintain temporal context across turns, (2) event memory architectures that preserve temporal relationships, (3) multi-turn temporal consistency frameworks that ensure coherent temporal reasoning throughout dialogues, and (4) temporal reference resolution systems that can handle context-dependent temporal expressions reliably.
F.3 Diachronic and Synchronic Knowledge Integration
Temporal questions often require combining diachronic sources (which capture change over time) and synchronic sources (which reflect knowledge at a specific moment). Most TQA systems treat these separately Wang et al. (2022); Gruber et al. (2024); Piryani et al. (2024b), limiting their ability to fully answer questions that span both historical context and present-day facts.
For instance, the question “How has unemployment changed since 2008, and what is the current rate?” requires integrating long-term trends from news archives (diachronic) with recent statistical summaries (synchronic). Current models often handle only one type of source effectively.
This integration requires more than merging sources; it demands alignment across different temporal granularities and anchoring schemes. Conflicts between historical accounts and retrospective summaries can further complicate the reasoning process.
Future systems should develop: (1) temporal alignment algorithms that can map events across different temporal representations, (2) cross-source reasoning frameworks that can weight and combine evidence from different temporal paradigms, (3) conflict resolution mechanisms for handling disagreements between diachronic and synchronic accounts, and (4) unified temporal representations that can accommodate both evolving and static knowledge seamlessly.
F.4 Temporal Uncertainty and Confidence Modeling
Many historical events involve uncertain or approximate dates, yet most TQA systems treat all temporal information as exact. This creates a mismatch with real-world complexity, particularly in multi-step reasoning where uncertainty can propagate and amplify.
For instance, the question "What happened between the fall of Rome and the beginning of the Renaissance?" involves fuzzy boundaries: the fall of Rome (often dated to 476 AD) and the Renaissance (spanning the 14th to 16th centuries) are debated and regionally variable. Systems must represent and reason over this imprecision, not resolve it artificially.
Temporal uncertainty also arises in natural language expressions like “around that time” or “in the late period,” and from conflicting historical accounts. Ignoring these leads to overconfident and potentially misleading answers.
Future systems should incorporate: (1) probabilistic temporal representations that can model uncertain dates and durations, (2) confidence propagation mechanisms for multi-step temporal reasoning, (3) uncertainty-aware answer generation that communicates temporal confidence to users, (4) conflict-aware reasoning that can handle and explain disagreements between sources, and (5) evaluation frameworks that reward appropriate uncertainty rather than false precision (Vashishtha et al., 2019; Zhou et al., 2021; Fatemi et al., 2024).
F.5 Multilingual and Multimodal Temporal QA
Temporal expressions, date formats, and cultural references vary widely across languages and modalities, yet most TQA systems are developed primarily for English and textual input. This poses challenges in multilingual and multimodal settings where temporal cues appear in varied forms such as non-Gregorian date formats in documents, seasonal imagery in videos, or handwritten timestamps in scanned texts. Current systems struggle to interpret these diverse signals due to limited cultural grounding and inadequate multimodal integration. For example, a model may fail to recognize a lunar calendar reference in Arabic or temporal context from a video of a snowstorm, suggesting winter.
Recent benchmarks like TimeDial Qin et al. (2021) and TOMATO Shangguan et al. (2024) highlight the need for culturally and visually grounded temporal reasoning. Future research should focus on developing multilingual temporal taggers, temporally annotated datasets in low-resource languages, and cross-modal alignment techniques that jointly reason over text, images, and video to capture time-related meaning in diverse cultural contexts Qin et al. (2021); Shangguan et al. (2024); Wu et al. (2025).
F.6 Implicit Temporal Intent Understanding
Many temporal questions conceal their intended timeframes, making implicit temporal intent one of the most difficult challenges in TQA. For example, the question, "What caused the economic crisis during Trump’s presidency?" could refer to events between 2017-2021 (e.g., COVID-19 or trade-related downturns) or more recent crises if the question is asked retrospectively. Current systems often fail to detect such ambiguity, either assuming the most recent crisis, misinterpreting the temporal anchor, or overlooking the ambiguity entirely.
This challenge stems from various factors: (1) conversational context, where prior dialogue establishes implicit timeframes; (2) cultural assumptions, such as shared historical reference points; (3) domain conventions, where temporal scope is understood implicitly; and (4) individual user perspectives or real-time situational context (e.g., ongoing events).
Existing models rely heavily on explicit temporal expressions and struggle with contextual inference (Zhou et al., 2021; Allein et al., 2023; Chu et al., 2024). Future work should focus on: (1) contextual intent detection models, (2) user and domain-aware temporal reasoning, (3) clarification strategies for resolving ambiguity, and (4) evaluation frameworks for temporal intent detection across diverse scenarios.
F.7 Evaluation and Benchmarking for Temporal Reasoning
Current evaluation practices for TQA largely rely on standard IR/NLP metrics such as accuracy, F1, MRR, or NDCG. While useful, these metrics often overlook the specific demands of temporal reasoning, such as correct temporal anchoring, event ordering, consistency across multi-hop inferences, and robustness to temporal ambiguity.
In Temporal QA, benchmarks like TimeQA (Chen et al., 2021a), MenatQA (Wei et al., 2023), ComplexTempQA Gruber et al. (2024) and TimeBench (Chu et al., 2024) provide test sets that target temporally sensitive tasks, such as temporal entailment, counterfactual reasoning, and timeline ordering. However, evaluation often reduces to span-level accuracy or multiple-choice correctness.
In TIR, metrics sometimes integrate recency weighting or time decay functions (Berberich et al., 2010), but still conflate temporal and topical relevance.
Future work should aim to develop: (1) evaluation metrics that capture temporal grounding and coherence, (2) metrics sensitive to ambiguous or conflicting time signals, and (3) unified protocols for comparing temporal capabilities across models and tasks.
| Method | Type | Task | Temporal Representation | Temporal Signals | Architecture |
| TempoT5 Dhingra et al. (2022) | Encoder-Decoder | TQA | Explicit timestamps | Document timestamps + temporal context | T5 + temporal conditioning |
| TempoBERT Rosin et al. (2022) | Encoder | Semantic Change Detection, Sentence Time PredictionSemantic Change Detection, Sentence Time Prediction | Explicit timestamps | Time tokens | BERT + temporal masking |
| BiTimeBERT Wang et al. (2023) | Encoder | Event Time Estimation, Document Dating, TQA | Explicit timestamps and content temporal expressions | Document timestamps + temporal expressions in text | BERT + dual temporal encoding |
| TALM Ren et al. (2023) | Temporal adaptation + Hierarchical modeling | Historical Dating | Implicit (Time-specific word variants) | Time-specific representations | BERT + Temporal Adaptation + Hierarchical Document Encoder |
| SG-TLM Su et al. (2023) | TLM | Temporal Language Modeling, TQA | Implicit Temporal Information | Syntactic role distributions + timestamps | BERT + syntax-guided masking + temporal-aware masking |
| TSM Cole et al. (2023) | Encoder | Temporal language modelling, TQA | Implicit temporal spans | Temporal expressions (dates, durations) | T5 + temporal span masking |
| Temporal Attention Rosin and Radinsky (2022) | Encoder | Semantic Change Detection | Explicit Document timestamp | Document timestamp | Transformer encoder with time-aware attention via time matrix T |
| TempRetriever Abdallah et al. (2025) | Neural Embedding | TQA | Explict+Implicit temporal information | Query timestamps + document timestamps | Semantic + Temporal Encoder |
| TsContriever Wu et al. (2024) | Neural Retrieval | TQA | Explict+Implicit temporal information | Query timestamps + document timestamps | Semantic + Temporal Encoder |
| TempRALM Gade and Jetcheva (2024) | RAG | TQA | Explicit temporal Information | Query + document timestamps | RAG + temporal enhancement |
| TimeR4 Qian et al. (2024) | RAG | TQA | Explicit temporal facts | Question constraints + TKG timestamps | Retrieve-Rewrite-Retrieve-Rerank with contrastive time-aware retriever + LLaMA2 reasoning |
| MRAG Siyue et al. (2024) | Neural + Symbolic | TQA | Explicit temporal information | Query constraints + Document timestamps | Modular RAG framework ( Question Processing, Retrieval+ Summarization +Semantic-Temporal Hybrid Ranking |
| Method | Input | Output | Knowledge | Processing | Methodology |
| TempoT5 Dhingra et al. (2022) | Text + temporal prefix | Masked span prediction | CUSTOMNEWS + TEMPLAMA | Temporal conditioning + uniform sampling | Input prefixing for time-aware modeling and efficient updates |
| TempoBERT Rosin et al. (2022) | Text + timestamps | Time tokens | News corpus | Implicit temporal reasoning via time masking | Time-aware pretraining with masking |
| BiTimeBERT Wang et al. (2023) | News articles with timestamps | Time predictions (day/month/year granularity) | NYT corpus | Bi-temporal reasoning using both timestamp and content time | Time-Aware Masked Language Modeling, Document Dating |
| TALM Ren et al. (2023) | Historical text | Time period classification | Chinese Twenty-Four Histories + English Royal Society Corpus | Temporal word adaptation and contextual learning | Learn separate word representations per time period + alignment |
| SG-TLM Su et al. (2023) | Text + timestamp | Masked token + time prediction | WMT News Crawl + Reddit Time Corpus | Syntax-guided masking (SGM) + temporal-aware masking (TAM) | Syntax-based lexicon selection for efficient temporal adaptation |
| TSM Cole et al. (2023) | Text with temporal expressions | Masked span prediction | Wikipedia | Temporal span masking using SUTIME parser + intermediate training | SUTIME parser identifies temporal expressions (Time, Duration, Set, Date) for targeted masking during intermediate pretraining |
| Temporal Attention Rosin and Radinsky (2022) | Text + Document timestamp | Time-specific contextualized embeddings | Time-annotated corpora (SemEval datasets) | Time-aware self-attention mechanism with additional time matrix | Extends self-attention with time matrix T to condition attention weights on temporal context via time-aware dot products. |
| TempRetriever Abdallah et al. (2025) | Query + Query timestamp + Document + Document timestamp | passages with temporal +semantic relevance | NYT + ChroniclingAmerica Corpus | Temporal alignment + Temporal understanding | Learned temporal embeddings with fusion techniques + Time-based negative sampling |
| TsContriever Wu et al. (2024) | Time-sensitive questions | Top-k temporally relevant documents | Nobel Prize benchmark corpus + Wikidata QA pairs | Contrastive learning + Query-side fine-tuning + Query routing | Supervised contrastive learning with time-aware query tuning and routing |
| TempRALM Gade and Jetcheva (2024) | Time-sensitive queries | Answer | Knowledge base | Temporal proximity-based document ranking | Temporal scoring |
| TimeR4 Qian et al. (2024) | Time-sensitive queries | Answer (entity or timestamp) | Knowledge base | Retrieve-Rewrite-Retrieve-Rerank pipeline with contrastive learning and LLM finetuning | Temporal fact retrieval, question rewriting, time-aware reranking, and LLM finetuning |
| MRAG Siyue et al. (2024) | Temporal questions + Document | Ranked passages | Wikipedia | Symbolic temporal scoring + Semantic-temporal hybrid ranking | Question decomposition (MC + TC) + Symbolic temporal |